A Complete Bibliography of Publications in IEEE Concurrency

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 2.07

Title word cross-reference

109-bit [Pri00]. 1200 [Chw98].

2 [ALPP00]. 2.0 [Chw98]. 2nd [YM97].

3.0 [CSFS00]. 30 [Cla98a].

95 [CS98]. ’97 [Sri98]. ’98 [MS99].

Abstractions [FP97]. academia [HR97]. acceptance [TBS98]. Access [MFGH99, Ano98g]. Accurate [KS98]. Achieving [Mil00a]. act [Chw98]. Active [GB99]. ActiveX [Chw98]. Actors [BN99, Bri98, KB98]. Ada [CS98]. Add [CKS99]. Advanced [Vaz98]. Advances [ALPP00]. Agent [Agh98d, KT98, Ken00, Mil00b, RRR^{+}99, SHG00, Mil99b]. Agent-Based [RRR^{+}99]. Agents [Bri98, GB99, KB98, MGR^{+}00, TR00, WS00, dRPT98]. ahead [CSFS00]. Aided [AKWS00]. air [RBHG00]. AL [DCH^{+}98]. Alamos [CKS99]. Algebra [ALPP00, Wyr00]. Algorithm [GZ98]. Algorithms [ALPP00, Mar98, Yal99a]. Alligator [LAS97]. American [Ano97m]. Amit [Jaj98]. Analysis [AMMV98, Hun00, Ken00, KK97, Zal97c]. anatomy [Pas99]. anger [Cha99]. Annual [Ano97b, Ano00a]. Apples [CST99]. Application [DJ98, HSSO97, Jos97, Kri97b].

Application-centric [Kri97b]. Applications [Cal98, CLZ99, FP97, Gom97, GGM97, Jos97, Kri97b, Kri97c, Kri97d, Kri98a, Kri98b, Kri98c, Kri98d, Kri99a, Kri99b, Kri99d, Kri99c, Kri00a, Kri00b, Lea98, MFGH99, Mil00b, PES^{+}98, WC98, Ano98g, Bac99d, Mil99b, Van98, Wei98]. Applied [Mil00c]. Approach [BN99, Gar99, Zal97b, Stp97, Vaz98]. Approaches [BH98]. Architecture [Gar99, MRF^{+}00, RRR^{+}99, Vaz98]. Architectures [Kra97, PTT98, Agh98b, Kri97c, Pas98]. Area [Stp97, vSHT99]. arena [Kri97a]. Aristotle [MS98b]. Armadillo [LAS97]. Array [SPF98, SPF99]. Art [Lev98]. Article [Ano97c, Ano97d]. Artificial [Trz97]. ASCI [Cla98a]. Aspects [Dev98]. asset [GS00]. Assignment [HA98, KA98]. Association [Zak99]. Assurance [AJ99]. Asynchronous [LBB97]. ATM [HDM^{+}97]. Atomic [Tha00]. automated [RBHG00]. Automatic [DK00]. Autonomous [ALPP00, GB99]. Avoidance [GZ98]. Award [Ano97m, Jaj99b].

Back [WSFC99, FG00a]. Balancing [Bes97, BN97, CLZ99, HC99]. Based [Bes97, Jos97, KC00, Mil00a, RRR^{+}99, Zal97c, AJ97b, Sat99]. Beauty [MS98b]. Behavior [Lu00]. Benchmarking [Mac97]. Better [LAS97]. between [CPSG00]. beyond [Cla98a, Sri98]. big [Chw98]. Billions [WSFC99]. bit [Pri00]. blends [Chw98]. blue [Cla97c, CSFS00]. Bluer [Cla97c]. Book [ALPP00, Ani97, Cob98, Dun99a, Dun99b, GS00, Jos97, Lev98, Mar98, Mik97, Möl98b, Möl98a, Pal99, Pap97a, Pap97b, Pap00, Pas98, Pas99, Pat00, Per97, Ran97, RI99, Sta98, Sta99, Stp97, Stp98, Szm97, Szy98, Tah98, Tro98, Trz97, Uma98a, Uma98b, Vaz98, Wei98, Wyr00, YM97, Yal98, Yal99a, Yal99b, Zal97a, Zal97b, Zal97c, d’A98, d’A99, vdV98].

Books [Ano97j, Ano97k, Ano97l, Ano98f, Cob98, Zal97b, Jos97, YM97, Zal97c]. Bottlenecks [HSSO97]. brain [EFH^{+}00]. brave [Bac99e]. bricks [Cla97a]. Bridging [HR97]. Broadcast [Asi98, Ber98, EFH^{+}00, HR97, Kah97, LK99, Sat99, Stü97, Van98]. broader [Ano97i]. Brokering [Elv97]. Building [Jos97, Mik97, TJD98, Wei98]. Bus [Mil00a, GPPR97]. Bus-Based [Mil00a].

C [Chw97, dD00]. Cache [SP00, THL00]. Caches [MIB00]. Caching [Bes97, CPSG00, HHT00, Mil00f]. CAIDA [Pri00]. Calendar [Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d]. Calendar/Call [Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d]. Caliper [Hun00]. calls [Wal98]. capacity [CSFS00]. Card [HH99, Hus99]. Carte [Bur97]. CASCH [AKWS00]. Case [MLB^{+}97, WA00]. Casting [PTT98]. Center [Kah97, Ano98g]. centric [Kri97b]. Challenges [GJ97, Kri99a, Sch97, vdV98, Bac00b, Hed97b]. changing [Ahm00d, Lew99]. Checkpointing [HJC^{+}99]. Chief [Agh97a, Agh97b, Agh97c, Agh98d, Agh98b, Agh98c, Bac99e, Bac99b, Bac99c, Bac99d, Bac00a, Bac00b, Bac00c, Bac00d]. Chip [KK97]. Christmas [Hed98]. classroom [Wei00]. Click [WSFC99]. Client [Jos97, Stp97, Uma98b]. Client/Server [Jos97, Stp97]. Client/Server [Uma98b]. climate [Lew99]. closing [Bac00d]. Cluster [Ahm00d, Ahm00a, Ahm00c, Ahm00b, CKS99, Tay97, Pat00, Sat99]. Clusters [Bac99b, HJC^{+}99, Ahm00c, Ahm00b, CST99, PKC97, Dun99a]. CoCa [vdSAW99]. Coherence [FMST00, WRdS00]. Coins [WSFC99]. collaborate [Ano97m]. Collection [ALPP00]. Collision [GZ98]. Collision-Avoidance [GZ98]. COM [WC98]. commerce [Hed98, Men00]. Commit [Tha00]. Committee [CKS99]. Communication [Pal99, Kri00a, PKC97]. community [Bac00a, CKS99]. Comparisons [LAS97, DCH^{+}98]. COMPaS [Sat99]. compiler [Chw97]. Complete [Per97, YM97, RS97]. Completeness [RI99]. Complex [ERH97, RWMB98, CKS99]. Complexity [Pal99]. Component [KC00]. Component-Based [KC00]. Compositional [Agh98b]. Comprehension [DK00]. comprehensive [ALPP00]. Computation [ALPP00, RI99, Szy98]. Computational [ALPP00, Tah98, VS99]. Computations [CC97]. Computer [AKWS00, AJ99, ERH97, Gar99, Pas98, Vaz98, Bac99c]. Computer-Aided [AKWS00]. Computers [Pap97b, Ahm00d, ALPP00]. Computerworld [Ano97m]. Computing [Ahm00d, Ahm00a, Ahm00c, Ahm00b, Ani97, DJ98, HDM^{+}97, HTS98, JC98, Jon99, KA98, LMR^{+}97, MS98b, MS99, Pot99, Ran97, Sri98, Tha00, ZEGF99, d’A98, vdV98, Agh98c, CKS99, Hed97b, Jos97, Kri97a, Kri99b, LK99, MS98a, Pas99, Stü97, YZ99, d’A99, Kah97, GS00, Lev98, Pal99, Pap97a, Stp98, Uma98a, Uma98b, Pap00, Pat00]. Concepts [Mik97, Möl98a, Bac99d]. Concurrency [CS98, MLB^{+}97, Szm97, Agh97c, Ano97i, Ano98g, Bac00d]. Concurrent [Agh97a, KS98, Möl98b, Sch97, Sta98, Tha00, UHS97, Yal98, Agh98c, Bri98, Lea98]. Conference [Stp98]. Conferences [TBS98]. Configurability [FMST00]. conflicting [AJ97b]. Connections [HDM^{+}97]. Considerations [Kri00a, PTT98, Kri99c]. Constructing [Kra97, SS98]. construction [CKS99]. content [Kri99a]. Contention [BN97, Dan99]. continue [Agh98d]. continues [CKS99]. Continuous [HHT00]. contract [CKS99]. Control [DS00, Cal98, RBHG00]. Cooperative [MIB00]. Coordinated [Zal97b]. coordination [RBHG00]. Corba [Cob98, FG00b, LS99, THL00, Chw98, CSFS00]. core [CST99]. coverage [Ano97i]. cozy [Ano97m]. CP [Asi98]. CP-PACS [Asi98]. cracked [Pri00]. Craft [Chw97]. Craft-y [Chw97]. create [Ano98g]. Creating [KS98]. Cross [Agh98c]. Cross-fertilizing [Agh98c]. cryptographic [Pri00]. culmination [Kri97a]. Current [Dev00]. curriculum [Men00]. Customization [WC98]. Cyber [Cla97a]. Cyber-bricks [Cla97a].

D [Stü97]. DAME [CC97]. Data

[BH98, Chw98, CC97, Jaj99a, Lu00, SP00, Ski99, WA00, WSFC99, ZEGF99]. Data-Parallel [CC97, WA00]. Database [ALPP00]. Databases [AJ97b, AJ97a, AJ99, Jaj98, Jaj99a, Jaj99b]. datasets [CKS99]. Dealing [Jos97]. Debugging [WA00, Zal97c]. decade [Szm97]. Defense [CKS99]. Defining [Bac99c]. delivering [Kri99a]. Delta [WRdS00]. Democracy [Hed97a]. democratization [Lea98]. Department [CKS99]. Dependence [KC00]. dependent [Kri99b]. depends [MS98b]. Design [Gom97, KT98, Ken00, Vaz98, WS00, Zal97b, GS00, Möl98a, Tro98]. Designing [HJC^{+}99, Mik97, SL00, Men00]. Develop [UHS97]. Developing [RBHG00]. Development [FP97, Zal97c, CSFS00, Sta99]. Developments [Dev00]. device [Cla99]. Did [Ani97]. Differences [CPSG00]. Diffusive [CLZ99]. disciplines [Agh98c]. Disks [HHT00]. Dissemination [TR00]. Distributed [AJ97b, AJ97a, AJ99, Ani97, BN99, Cal98, CPSG00, Cob98, CC97, CE97, Dan98, DS00, ERH97, FG00a, FP97, Gar97, GS00, Gom97, GJ97, GGM97, HDM^{+}97, Jaj98, Jaj99a, Jaj99b, KA98, KC00, Kra97, Lu00, Mar98, MLB^{+}97, Mil00f, MGR^{+}00, Möl98a, PES^{+}98, Ran97, RWMB98, SL00, THL00, TR00, Trz97, Uma98a, Zak99, Zal97b, ZEGF99, vSHT99, Ano98g, Bac00a, Bac00b, Chw99, Hed97b, JG97, Jos97, Pap00, RS97, WC98, Zal97c]. doing [Ahm00c]. Don’t [Hus99]. Door [Ani97]. Driven [CE97, GPPR97]. Driving [SvdAA99]. DS [Bac00a, Bac00d]. DSM [FMST00]. DSP [KK97]. Dynamic [CLZ99, Dan98, HA98, HC99, RRR^{+}99, RBHG00]. Dynamics [Tah98].

e-Christmas [Hed98]. e-commerce [Men00]. Economy [SvdAA99]. Editor [Agh97a, Agh97b, Agh97c, Agh98d, Agh98b, Agh98c, Bac99e, Bac99b, Bac99c, Bac99d, Bac00a, Bac00b, Bac00c, Bac00d, Dev98, Dev00, Jon99, Mil00b, Mil00c, Mil00f, PYZ00]. Editor-in-Chief [Agh97a, Agh97b, Agh97c, Agh98d, Agh98b, Agh98c, Bac99e, Bac99b, Bac99c, Bac99d, Bac00a, Bac00b, Bac00c, Bac00d]. Editors [Ano97i, CM97, ERH97, GJ97, KB98, SW99, ZEGF99]. Education [BAB98, Bis00, FG00a, Men00, Tri99, Wei00]. effective [YZ99]. Effects [SHG00]. Efficiency [CC97]. Efficient [MIB00, PKC97]. effort [CKS99]. Eighth [Stp98]. electronic [Hed98]. Elements [GS00]. Eliminating [HSSO97]. Embedded [CSFS00, DS00, HR97, HSSO97, MLS^{+}00]. emerging [Agh97b]. enable [RBHG00]. Encryption [WSFC99]. end [WSFC99]. Energy [vdSAW99]. Engineering [ERH97, GJ97, Jos97, LMR^{+}97, Agh97b, JG97, Mik97]. engineers [Pri00]. Enhanced [dRPT98, EFH^{+}00]. Enhancing [AHA^{+}00, Kri00b]. Enterprise [Bus99]. Enterprise-Wide [Bus99]. Environment [CC97, LLV98]. Environments [RWMB98, Kri99a, Jos97]. Europe [Hed98].

European [Ber98, GCSS97, Stü97]. Europort [Stü97]. Europort-D [Stü97]. Evaluation [Dan98, LS99]. events [Ahm00a]. evolution [Sch99]. Evolving [Bac99d]. Exception [dD00]. Executions [KS98]. Expectations [Bac00b]. experts [Ano97m]. Extending [HDM^{+}97]. Extensible [KP98, SS98]. extensions [JC98]. Extreme [BAB98].

face [Ahm00d]. Facility [Chw98]. fact [TBS98]. Factory [GPPR97]. failures [RS97]. fall [CKS99]. fantasy [TBS98]. fashion [CKS99]. Fast [MRF^{+}00, PKC97]. Fault [AHA^{+}00, AJ99, FMST00]. FC [DCH^{+}98]. FC-AL [DCH^{+}98]. feature [Ano97m]. Features [CS98]. fertilizing [Agh98c]. Field [Ano97m]. File [KP98, SHG00]. Filtering [TR00]. final [Bac00c]. flawless [GS00, Pap00]. Flexible [Lu00]. Fluid [Tah98]. flypaper [Chw98]. Focus [Cha99, Cla97a, Cla98a, Cla98b, Lew99]. Formal [Zal97b]. Fortran [Lev98]. forum [Kri97d]. forward [FG00a]. Foundations [Trz97]. Framework [Hun00, Kra97]. Free [Chw97]. Frigate [KP98]. frontier [LK99]. functional [EFH^{+}00]. Fundamentals [Cob98]. Future [Mil99c, vdV98, TBS98].

gaining [Jaj99a]. Garf [GGM97]. Gasser [Bri98]. Gene [CSFS00]. general [TBS98]. general-purpose [TBS98]. Generating [GPPR97]. Generation [Uma98b, Kri00a, Sri98]. Genetic [Ahm00b]. German [Ano97m]. get [Ano97m]. Getting [Chw97]. Gigantic [Ahm00c]. glance [Ahm00a]. Global [Asi98, Ber98, EFH^{+}00, Gar97, HR97, Kah97, LK99, Sat99, Sta97, Stü97, Van98]. Globe [vSHT99]. Go [Chw99, Jon99]. good [Cla97b]. grand [Bac00d]. graphic [Chw97]. Gray [Jaj99b]. grid [LK99, CKS99]. gridlock [CST99]. Grids [Wei00]. Groups [FG00b]. grow [Agh98d]. Growing [CST99]. Guest [CM97, Dev98, Dev00, ERH97, GJ97, Jon99, KB98, Mil00b, Mil00c, Mil00f, SW99, ZEGF99]. Guide [WSFC99].

Handbook [Ran97]. handle [RS97]. Handling [LBB97, dD00]. Hardware [Gar99]. Hardware/Software [Gar99]. harnessing [CM97]. Hector [RRR^{+}99]. helps [Pri00]. Heterogeneous [AMMV98, KA98, ALPP00]. Hierarchical [HHT00, HJC^{+}99]. Hierarchy [PTT98]. High [ALPP00, Bur97, Kah97, Kri98a, Mil00a, Pap97a, Pat00, TJD98, VS99, vdSAW99, vdV98]. High-Energy [vdSAW99]. High-Performance [ALPP00, Bur97, Kah97, VS99, Kri98a, Pap97a]. High-Speed [TJD98]. Highway [HTS98]. HIPPI [HDM^{+}97]. Hive [MGR^{+}00]. Home [Hus99, Ano98g]. Hot [Dan99]. Hot-Spot [Dan99]. House [Hed97a]. HP [Ano97m, Hun00]. HPC [Ano97m, Van98]. HPCN [Ber98, GCSS97]. HPF [Chw97]. Huge [SPF98, SPF99]. human [EFH^{+}00]. hunger [WSFC99]. Hype [BBE^{+}99]. Hypersequential [UHS97].

I/O [HJC^{+}99, MFGH99]. I/O-Intensive [MFGH99]. IBM [Cla97c]. I’d [WSFC99]. IEEE [CKS99]. II [MRF^{+}00]. Image [MFGH99]. Implementation [CGS^{+}97, Ken00]. Implementing [Lu00]. Improve [HTS98]. Improved [HC99]. Index [Ano97b, Ano98e, Ano99e, Ano00a]. industrial [Van98]. Industry [Cla97b, Hed98, HR97, Stü97]. Information [LK99, TR00, Bis00, Kri98b, Tri99, WSFC99]. Infoverse [Kri99d]. Infrastructure [WS00]. innovation [Cha99]. integer [Ano97m]. integer-sorting [Ano97m]. Integrated [RWMB98]. Integrating [BH98]. integrity [AJ97a]. Intel [Ano97m, BN97]. Intelligence [Trz97]. intelligent [Cla99, Hed97b]. Intensive [MFGH99].

Interacting [Zal97b]. Interactive [PES^{+}98]. Interconnection [LAS97]. Interconnections [d’A98, d’A99]. Interface [DCH^{+}98]. interleavings [BAB98]. Internet [Jos97, Cha99, CST99, Kri98a, Mil00d, dRPT98]. interoperability [Bac99b]. Interview [Bri98, Cla97c, Jaj98]. Introduction [CM97, Dev98, Dev00, ERH97, GJ97, Jon99, KB98, Mar98, Mil00b, Mil00c, Mil00f, Pap97a, SW99, Tah98, ZEGF99]. Invocation [Wal98]. IPA [HR97]. IPC [MLB^{+}97]. Irix [Chw98]. issue [Bac00c]. Issues [KT98]. iWarp [Pas99].

Japanese [Asi98]. Java

[GS00, Pap00, Yal98, Agh97a, Chw98, CSFS00, DS00, Wal98]. JavaCard [BBE^{+}99]. Jim [Jaj99b].

Keeping [Tri99]. Kernel [DS00]. key [Pri00]. Knowledge [Zal97c]. Knowledge-Based [Zal97c].

lab [CST99]. Laboratories [PES^{+}98]. Laboratory [CKS99]. Labview [Chw98]. Lancaster [WSFC99]. Language [WS00, Yal98]. large [Bac00b, TBS98]. large-scale [Bac00b, TBS98]. Layout [Tro98]. leads [Chw98]. learning [FG00a, RBHG00]. Leave [Hus99]. Legacies [Jos97]. Letter [PYZ00]. library [Chw97]. lifestyle [Chw98]. like [WSFC99]. Limits [RI99]. line [CKS99]. Linear [ALPP00, Wyr00]. Linux [Pri00]. Load [Bes97, CLZ99, Dan98, HC99]. Load-Balancing [CLZ99]. Local [Stp97]. Location [Kri99b]. Location-dependent [Kri99b]. Lock [Ani97]. look [Asi98, ALPP00]. Loop [HKS^{+}97]. LosLobos [Pri00]. Low [Pot99]. Lung [Ano97m].

Mach [MLB^{+}97]. Machines [Sta97, TBS98]. Magazines [Bac99c]. Maintenance [HTS98]. Malaysia [WSFC99]. managed [Cla97c]. Management [AHA^{+}00, ALPP00, Bus99, KC00, RRR^{+}99, SW99, THL00, RBHG00]. Managing [Kri98b]. Manipulators [GZ98]. market [Hed98]. marks [Bac00d]. masses [Cla98b]. Massively [HTS98]. Matching [Dec97]. Mathematics [ALPP00]. matrices [ALPP00]. Matrix [ALPP00]. Media [HHT00]. Meeting [Sch97]. meetings [Hed97a]. Memory [CPSG00, Dan99, JM99, Mil00a, PTT98, Ano97m]. Memory-Hierarchy [PTT98]. Message [LBB97]. Messages [PKC97]. Method [Wal98]. Methods [Gar97, Gom97, Tro98, Szy98]. Microphone [SPF98, SPF99]. Middleware [Mil99a]. Might [HTS98]. Migration [Chw98]. millenium [Jaj99b]. millennium [Bac99e]. Mining [CKS99, Ski99, Zak99, ZEGF99]. MiSFIT [SS98]. MobiCom [MS99, Sri98]. Mobile [JC98, Jon99, KT98, Kri99d, Kri99c, MS98b, MS98a, MS99, Mil99b, Pot99, SHG00, Sri98, TR00, WS00, Kri99a, Kri00a, Kri00b]. Mobile-Agent [KT98]. mobility [Bac99b]. Model [CE97, vdSAW99, Men00]. Model-Driven [CE97]. Modeling [Ken00, Lew99]. Models [LBB97, Szy98]. Modular [BN99, LLV98]. money [CKS99]. Monitoring [Zal97c]. moves [CSFS00]. MPI [Per97, YM97, YM97, YM97]. MPPs [PKC97]. MRI [EFH^{+}00]. MSC [WSFC99]. Multicasting [TJD98]. multidatabases [ALPP00]. Multimedia [Kri97b, Kri97a, Kri97c, Kri97d, Kri98a, Kri98b, Kri98c, Kri98d, Kri99a, Kri99b, Kri99d, Kri99c, Kri00a, Kri00b, Kri97d]. Multiparty [Zal97b]. multiple [RS97]. Multiprocessor [CM97, Dan99, KK97, Mac97]. Multiprocessors [Mil00a, GPPR97].

Naming [LS99]. Naming-Service [LS99]. National [CKS99, Kah97]. NCSA [Ano98g]. netcasts [Hed97a]. Nets [BN99]. Network [Ahm00d, Cla99, GS00, LAS97, TJD98, Wei98, Pri00]. Networked [SvdAA99]. Networking [MPC^{+}00, MGR^{+}00]. Networks [HDM^{+}97, Stp97]. News [Ano97m, Ano98g, Chw99, CST99, CKS99, CSFS00, Pri00, RBHG00, WSFC99, Cla97b]. Niño [Ano98g]. nirvana [Cla99]. No [Cla97b]. Nomadic [WS00]. nuclear [Cla97b]. nukes [Cla97b]. Numerical [ALPP00, Lev98].

O [HJC^{+}99]. O-Intensive [MFGH99]. Object [Dev98, Dev00, FG00b, Jos97, Lea98, THL00, WC98]. Object-Oriented [Dev98, Dev00, WC98]. objectives [AJ97b]. Objects [Cob98, GB99, Bri98]. Observing [Gar97]. offers [ALPP00]. older [Chw98]. one [FG00a]. online [Hed97a, Bac00a, Bac00d]. OO [KP98]. open [Bac99b]. OpenCard [HH99]. opening [Bac00d]. OpenMP [Cla98b]. Operating [Mil99c, Mil00c, Möl98a, CM97, CGS^{+}97]. Opportunities [GJ97]. Optical [d’A98, Chw97, d’A99]. Optimal [KA98]. Optimize [POS^{+}00]. Optimizing [TR00]. Orbix [Cob98]. ORBs [POS^{+}00]. Oriented [Dev98, Dev00, Jos97, Lea98, WC98]. Our [Bac99c]. outshines [Chw98]. Overview [JM99, ALPP00]. own [CST99].

P [LLV98, RI99]. P-Completeness [RI99]. P-RIO [LLV98]. Package [Wyr00]. PACS [Asi98]. page [Ano98g]. Papers [Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d]. Para [Ano98g]. Para-Scope [Ano98g]. paradigms [Cla97a]. Paragon [BN97]. Parallel [CKS99, CC97, DJ98, Dun99b, Gar99, GZ98, GJ97, HSSO97, HTS98, HKS^{+}97, Kri97d, Kri98c, Lev98, LLV98, LMR^{+}97, Mik97, MRF^{+}00, Pal99, Pap97b, Pas98, Ran97, RI99, Ski99, Stp98, Szy98, Tah98, Tro98, Uma98a, WA00, Wyr00, YM97, Yal99a, Yal99b, Zak99, ZEGF99, d’A98, d’A99, Cla98b, CGS^{+}97, JG97, Kri97b, Kri97a, Kri98b, LK99, Pas99, Sta99, Stü97, TBS98, YZ99]. Parallel-Loop [HKS^{+}97]. Parallel-Programming [LLV98]. Parallelism [BH98]. Parallelization [DK00, vdSAW99]. Parallelizing [MFGH99]. Part [SPF99]. Pathforward [Cla98a]. Patterns [Lea98, POS^{+}00]. PC [Sat99]. PC-based [Sat99]. Peace [CGS^{+}97]. Call [Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d]. Server [Jos97, Stp97]. Software [Gar99]. Performance [ALPP00, AMMV98, Bur97, HSSO97, Hun00, Kah97, KK97, Mil00a, VS99, YZ99, vdV98, CGS^{+}97, Kri98a, Pap97a, Pat00]. performs [Chw98]. Perspective [GCSS97]. perspectives [Szm97]. petaflops [CSFS00]. Petri [BN99]. Physics [vdSAW99]. Pict [WS00]. Pipelined [Pas98]. Plapack [Wyr00]. playing [Hed97b]. Policies [CLZ99]. popularity [Jaj99a]. Portable [MRF^{+}00, PKC97]. portal [RBHG00]. Portrayal [KS98]. ports [Chw98]. poses [Hed97b]. power [CM97, LK99, Pot99]. PQE2000 [Van98]. Practical [CS98, Dev98, Yal99b]. Practice [Pap97b]. Prediction [AMMV98]. Prefetching [HHT00]. Preserving [CC97]. Principle [POS^{+}00]. printing [YM97]. privacy [Mil00e]. Probability [Zal97b]. problem [Ano98g]. Problems [ALPP00, LAS97]. procedure [Wal98]. Proceedings [Stp98]. Process [DS00, RS97]. Processes [SvdAA99, Zal97b]. Processing [Kri97d, MFGH99, Stp98, AJ97b, Chw97, JC98, Kri97c, CKS99]. processors [CKS99]. Products [Chw97, Chw98]. Profit [YZ99]. Profit-effective [YZ99]. profitup [PYZ00]. Program [DK00, HKS^{+}97]. Programming [Dev98, FG00b, GGM97, KT98, LLV98, UHS97, YM97, Yal98, Ahm00b, Lea98, Cob98, Möl98b, Yal99b, Zal97a]. Programs [Mik97, UHS97, WA00, Sta98, Sta99]. Progress [Mac97]. Properties [Gar97]. Protocol [GS00]. Protocols [FMST00, WRdS00]. Proxy [CPSG00]. PS [AMMV98]. Pthreads [Zal97a]. purpose [TBS98]. Putting [CSFS00]. puzzle [WSFC99]. Pyl [Ber98].

QoS [Pot99, dRPT98]. quality [Kri00b].

race [CSFS00]. RAM [Chw97]. Ray [PTT98]. Re [Jos97]. Reactive [Zal97b]. Real [Chw97, HA98, POS^{+}00, Sta97, Zal97c]. Real-Time [HA98, POS^{+}00, Sta97, Chw97, Zal97c]. Reality [BBE^{+}99]. Reasoning [Kra97]. Recipes [Lev98]. record [Ano97m]. Recovery [RS97]. Reducing [Dan99]. Reduction [Bes97]. Referees [Ano97a]. Reference [YM97, Men00, Per97]. reflect [Bac00c]. reflections [Szm97]. Reflective [JM99]. Relative [WA00]. Reliable [GGM97, Wei98]. Remote [MLB^{+}97, Wal98]. Render [Elv97]. Rendering [PTT98]. replication [Jaj99a]. Requirements [Dec97]. research [ALPP00, Kri97d]. Resource [RRR^{+}99]. Rethinking [AJ97a]. Reuse [DJ98]. Review [Ani97, Dun99a, Dun99b, Lev98, Mar98, Mik97, Möl98b, Möl98a, Pal99, Pap97a, Pap97b, Pap00, Pas98, Pas99, Pat00, Per97, Ran97, RI99, Sta98, Sta99, Stp97, Stp98, Szm97, Szy98, Tah98, Tro98, Trz97, Uma98a, Uma98b, Vaz98, Wei98, Wyr00, Yal98, Yal99a, Yal99b, Zal97a, d’A98, d’A99, vdV98].

Reviewers [Agh98a, Bac99a, Bac00e, Bac00f]. Reviews [ALPP00, Cob98, GS00, Jos97, YM97, Zal97b, Zal97c]. revolution [Tri99]. RIO [LLV98]. RoboCup [Hed97b]. Robot [GZ98]. Robots [Hed97b]. Role [Ken00, Mil99a]. Roundtable [Bur97, Dec97, FP97, Gom97, JG97, Kra97, Sch97, Sta97]. running [Ano97m].

Safe [SS98]. Salute [Bac00e, Bac00f]. satisfying [AJ97b]. save [CKS99]. Scalable [Kri98d, MIB00, CKS99]. scale [Bac00b, TBS98]. Scheduling [AKWS00, HA98, HKS^{+}97]. Scientific [DJ98, Lev98, LMR^{+}97, Pap97a, Stp98, CKS99]. Scope [Bac99c, Ano98g]. Scoped [Lu00]. ScriptEase [Chw98]. SDL [SL00]. Search [Dun99a]. seasons [WSFC99]. Second [Uma98b]. Secure [Wei98]. Security [AJ99, Ani97, Mil00e, Bis00]. seeding [Hed98]. sees [Chw97]. Semantics [AJ97b]. Semantics-based [AJ97b]. Sempa [LMR^{+}97]. Sensitivity [Dan98]. Sequential [Sta98]. Server [Bes97, Uma98b]. Server-Based [Bes97]. Servers [HHT00, Kri98c, Kri98d]. Service [LS99, SHG00]. Services [Bur97, SL00]. Serving [Bac00a]. set [Hed97b]. sets [Ano97m]. shaping [Kri99d]. Shared [CPSG00, Dan99, Mil00a, PTT98, Ano97m, GPPR97]. shared-bus [GPPR97]. Shared-Memory [Dan99, Mil00a, PTT98, Ano97m]. Sharing [Dan98, Lu00]. Sheth [Jaj98]. Show [Tay97]. SIAM [Stp98]. siblings [Chw98]. simulates [CST99]. Simulation [BN99, Cla97b, GPPR97]. simulations [Ano97m]. Simulator [MRF^{+}00]. Single [HJC^{+}99, KK97]. Single-Chip [KK97]. size [MS98b]. Small [Chw98]. Smart [HHT00, HH99, Hus99, MS98a]. smarter [CKS99]. Smithsonian [Ano97m]. SMP [Sat99]. soccer [Hed97b]. Society [Bac99c].

Software [AJ99, DJ98, Gom97, GJ97, JG97, Kra97, LMR^{+}97, Mik97, RWMB98, Sch97, Zal97c, Agh97b, Agh98b, Kri97b]. solid [GS00, Pap00]. solve [WSFC99]. sorting [Ano97m]. Space [HJC^{+}99, Vaz98, Chw98]. spaces [MS98a]. spawns [Cha99]. Specification [DJ98]. Speed [TJD98]. Speedup [PYZ00]. Splitting [SP00]. Spot [Dan99]. Spotlight [Cal98, Cla97b, Cla99, Hed97b, Hed98, Wal98]. SSA [DCH^{+}98]. SSI [HJC^{+}99]. standard [Cla98b]. standards [Sch99]. State [Ano97m]. Steering [VS99]. steps [FG00a]. storage [Chw98]. Strategies [HC99, Ski99]. Structural [Kra97]. Structured [Sta99, ALPP00]. Study [MLB^{+}97, WA00]. Subword [JC98]. sufficiently [ALPP00]. Summaries [Ano97c, Ano97d]. supercluster [Pri00]. Supercomputer [EFH^{+}00]. Supercomputer-enhanced [EFH^{+}00]. supercomputers [Ano97m, Cla97b]. Supercomputing [CKS99, Asi98, Dun99b]. support [Ano98g]. Survey [SP00, Zak99]. Synchronization [BN97]. System [AMMV98, Ken00, KP98, MIB00, vSHT99, CGS^{+}97, WSFC99, Pas99]. Systems [AHA^{+}00, Bac99c, CC97, CE97, Dan98, Dan99, Dev00, DS00, ERH97, Gar97, GJ97, JM99, KA98, KT98, KC00, Kra97, Mil99c, Mil00b, Mil00c, MLS^{+}00, Mil00f, Möl98a, RWMB98, SW99, SS98, THL00, Zal97b, Ano97m, Bac00a, Bac00b, CM97, JG97, JC98, Kri98b, Kri00a, RS97, WC98, Zal97c, HR97, ALPP00, Zal97b, Zal97c, vdV98].

T3E [Chw98]. T3E-1200 [Chw98]. Taiwan [Kah97]. takes [CST99]. Talking [HH99]. talks [Ano97m]. tapestry [Agh97b]. Task [BH98, KA98]. Tasking [MLB^{+}97]. Tasks [HA98]. Technical [CKS99]. Techniques [HTS98, VS99]. technologies [Agh98d]. Technology [Jaj98, Mil00d, Tri99]. template [Chw97]. Templex [Chw97].

Temporal [Zal97b]. Tera [Ano97m]. teraflops [CSFS00, Ano97m]. test [CSFS00]. textbook [GS00]. Tflops [Cla98a]. their [CST99]. Theory [ALPP00, Pap97b, RI99]. these [CST99]. Thierry [Ber98]. Things [MGR^{+}00]. third [Sri98]. thought [Ano98g]. Three [Zal97b]. Time [BN99, HA98, POS^{+}00, Sta97, Zal97b, Bac00c, Chw97, RS97, Zal97c]. Timed [BN99]. today [Mil99a]. Tolerance [AHA^{+}00, AJ99, FMST00]. tomorrow [Mil99a]. Tool [AKWS00, GGM97, RWMB98]. Tools [DK00, GCSS97, Hun00, Mik97, RWMB98, RBHG00, Van98]. tourist [WSFC99]. town [Hed97a]. trace [GPPR97, GPPR97]. trace-driven [GPPR97]. Traffic [Bes97, RBHG00]. Trafficmaster [CST99]. train [Pri00]. transaction [AJ97b]. Trapper [HSSO97]. Tree [TJD98]. Trend [Mil99a, Mil99b, Mil99c, MLS^{+}00, Mil00d, MPC^{+}00, Mil00e]. trends [Ber98]. Tuning [HKS^{+}97, KK97]. Tunnel [MRF^{+}00, dRPT98]. Turing [Jaj99b]. Two [Cob98, DK00, FG00a, Jos97, YM97, Zal97c].

UCAR [Ano97m]. UK [Cha99]. Understanding [Dun99b]. University [RBHG00]. unmanned [Cal98]. updated [ALPP00]. Use [CS98]. User [Dec97, KP98]. User-Extensible [KP98]. Using [AMMV98, BN99, Cob98, HKS^{+}97, POS^{+}00, RS97, Wyr00, d’A98, WC98, d’A99].

vector [RS97]. vectors [CKS99]. vehicles [Cal98]. Verification [Sta98, Zal97b]. version [ALPP00]. versus [DCH^{+}98, PYZ00]. video [JC98]. Virtual [Bur97, Chw98, Dec97, FP97, Gom97, JG97, Kra97, Sch97, Sta97, RBHG00]. vision [Chw97]. visual [Kri00b]. Visualization [Elv97, HKS^{+}97, Zal97c]. VLSI [Tro98]. Volume [ALPP00, Ano98e, Ano99e, Ano00a, Elv97, PTT98]. Vote [CGS^{+}97]. Voyager [Chw98].

Warp [BN99]. Wars [Mil99a, Mil99b,

Mil99c, MLS^{+}00, Mil00d, MPC^{+}00, Mil00e]. was [Ano98g]. Way [UHS97, CSFS00]. weapons [Cla97b]. weather [Ano97m]. Web [Jos97, MIB00]. Web-Based [Jos97]. Welcome [Agh97c]. Where [Ahm00c]. White [Hed97a]. Wide [Bus99, vSHT99]. Wide-Area [vSHT99]. Widespread [TBS98]. win [Ano97m]. Wind [MRF^{+}00]. winner [Jaj99b]. wireless [WSFC99]. Wisconsin [MRF^{+}00]. Without [Hus99]. Workflow [AHA^{+}00, Bus99, Jaj98, SW99, Sch99]. Workflow-Management [SW99]. workloads [GPPR97]. Workshop [HR97, Kri97d]. Workstation [Tay97, PKC97]. world [Bac99e, WSFC99]. WWW [Bes97].

ZPL [WA00].

References

Agha:1997:ECC

[Agh97a] Gul Agha. From the Editor-in-Chief: Concurrent Java. IEEE Concurrency, 5(4):2–3, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                                  URL http://dlib.computer.
                                                  org/pd/books/pd1997/pdf/p4002.
                                                  pdf.

Agha:1997:ECE

[Agh97b] Gul Agha. From the Editorin-Chief: The emerging tapestry of software engineering. IEEE Concurrency, 5(3):2–4, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3002.
pdf.

Agha:1997:ECW

[Agh97c] Gul Agha. From the Editorin-Chief: Welcome to Concurrency! IEEE Concurrency, 5(1):2–4, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1002.pdf.

Agha:1998:R

[Agh98a] Gul Agha. 1997 reviewers. IEEE Concurrency, 6 (3):92–93, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3092.pdf.

Agha:1998:ECCa

[Agh98b] Gul Agha. From the Editorin-Chief: Compositional software architectures. IEEE Concurrency, 6(1):2–4, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1002.pdf.

Agha:1998:ECCb

[Agh98c] Gul Agha. From the Editorin-Chief: Cross-fertilizing concurrent computing disciplines. IEEE Concurrency, 6(3):2–4, July/September 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1998/pdf/p3002.
pdf.

Agha:1998:ECA

[Agh98d] Gul A. Agha. From the Editor-in-Chief: Agent technologies continue to grow. IEEE Concurrency, 6(4):2–4, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1998/pdf/p4002.
pdf.

Alonso:2000:EFT

[AHA^{+}00] Gustavo Alonso, Claus Hagen, Divyakant Agrawal, Amr El Abbadi, and C. Mohan. Enhancing the fault tolerance of workflow management systems. IEEE Concurrency, 8 (3):74–81, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3074.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3074abs.htm.

Ahmad:2000:CCGa

[Ahm00a] Ishfaq Ahmad. Cluster computing: a glance at recent events. IEEE Concurrency, 8 (1):67–69, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1067.pdf.

Ahmad:2000:CCGc

[Ahm00b] Ishfaq Ahmad. Cluster computing: Genetic programming in clusters. IEEE Concurrency, 8 (3):10–11, 13, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3010.pdf.

Ahmad:2000:CCGb

[Ahm00c] Ishfaq Ahmad. Cluster computing: Gigantic clusters: Where are they and what are they doing? IEEE Concurrency, 8(2):83–85, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2083.pdf.

Ahmad:2000:CC

[Ahm00d] Ishfaq Ahmad. Cluster computing: Network computers: The changing face of computing. IEEE Concurrency, 8(4):9–11, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4009.pdf.

Ammann:1997:DD

[AJ97a] Paul Ammann and Sushil Jajodia. Distributed databases: Rethinking integrity. IEEE Concurrency, 5(4):5–6, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p4005.
pdf.

Amman:1997:DDS

[AJ97b] Paul Ammann and Sushil Jajodia. Distributed databases: Semantics-based transaction processing: satisfying conflicting objectives. IEEE Concurrency, 5(2):8–10, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2008.pdf.

Ammann:1999:DDC

[AJ99] Paul Ammann and Sushil Jajodia. Distributed databases: Computer security, fault tolerance, and software assurance. IEEE Concurrency, 7(1):4–6, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1004.pdf.

Ahmad:2000:CTC

[AKWS00] Ishfaq Ahmad, Yu-Kwong Kwok, Min-You Wu, and Wei Shu. CASCH: a tool for computeraided scheduling. IEEE Concurrency, 8(4):21–33, October/ December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.

org/pd/books/pd2000/pdf/p4021.
pdf;   http://www.computer.
org/concurrency/pd2000/p4021abs.
htm.

Amestoy:2000:BRC

[ALPP00] Patrick Amestoy, Xiaoye Li, Fred Petry, and Marcin Paprzycki. Book reviews: Collection offers overview of research on structured matrices: High Performance Algorithms for Structured Matrix Problems, Advances in the Theory of Computation and Computational Mathematics, Volume 2; A comprehensive look at heterogeneous multidatabases: Management of Heterogeneous and Autonomous Database Systems; new version not sufficiently updated: Numerical Linear Algebra for High-Performance Computers. IEEE Concurrency, 8(4):91–93, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4091.pdf.

Aversa:1998:HSP

[AMMV98] Rocco Aversa, Antonino Mazzeo, Nicola Mazzocca, and Umberto Villano. Heterogeneous system performance prediction and analysis using PS. IEEE Concurrency, 6(3):20–29, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3020.pdf;   http:
//www.computer.org/concurrency/
pd1998/p3020abs.htm.

Anita:1997:BRS

[Ani97]  Sebastian Anita.  Book re-

view: Security in Distributed Computing: Did You Lock the Door? IEEE Concurrency, 5 (3):76–77, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p3073.pdf.

Anonymous:1997:R

[Ano97a] Anonymous. 1996 referees. IEEE Concurrency, 5(3):83, July/September 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/p3083.
                                        pdf.

Anonymous:1997:AI

[Ano97b] Anonymous. Annual index. IEEE Concurrency, 5(4):83– 86, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p4083.pdf.

Anonymous:1997:ASa

[Ano97c] Anonymous. Article summaries. IEEE Concurrency, 5 (3):3–4, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p3003.pdf.

Anonymous:1997:ASb

[Ano97d] Anonymous. Article summaries. IEEE Concurrency, 5 (4):3–4, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4003.pdf.

Anonymous:1997:CCPa

[Ano97e] Anonymous. Calendar/call for papers. IEEE Concurrency, 5 (1):91–97, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1091.pdf.

Anonymous:1997:CCPb

[Ano97f] Anonymous. Calendar/call for papers. IEEE Concurrency, 5(2):94–97, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2094.pdf.

Anonymous:1997:CCPc

[Ano97g] Anonymous. Calendar/call for papers. IEEE Concurrency, 5 (3):84–89, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3084.pdf.

Anonymous:1997:CCPd

[Ano97h] Anonymous. Calendar/call for papers. IEEE Concurrency, 5(4):87–89, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4087.pdf.

Anonymous:1997:ECB

[Ano97i] Anonymous. From the Editors: Concurrency’s broader coverage. IEEE Concurrency, 5 (2):2–3, April/June 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p2002.
pdf.

Anonymous:1997:NBa

[Ano97j] Anonymous. New books. IEEE Concurrency, 5(1):88, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1088.pdf.

Anonymous:1997:NBb

[Ano97k] Anonymous. New books. IEEE Concurrency, 5(2):90, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Anonymous:1997:NBc

[Ano97l] Anonymous. New books. IEEE Concurrency, 5(3):77, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3077.
pdf.

Anonymous:1997:NAG

[Ano97m] Anonymous. News: American, German supercomputers get cozy; lung and weather simulations win Computerworld Smithsonian Award; State of the Field talks to feature HPC experts; UCAR and HP to collaborate on shared-memory systems; Intel Teraflops is up and running; Tera sets integer-sorting record. IEEE Concurrency, 5 (3):78–80, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3078.pdf.

Anonymous:1998:CCPa

[Ano98a] Anonymous. Calendar/call for papers. IEEE Concurrency, 6 (1):95–97, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1095.pdf.

Anonymous:1998:CCPb

[Ano98b] Anonymous. Calendar/call for papers. IEEE Concurrency, 6(2):95–97, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2095.pdf.

Anonymous:1998:CCPc

[Ano98c] Anonymous. Calendar/call for papers. IEEE Concurrency, 6 (3):94–96, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p3094.pdf.

Anonymous:1998:CCPd

[Ano98d] Anonymous. Calendar/call for papers. IEEE Concurrency, 6(4):86–88, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p4086.pdf.

Anonymous:1998:IV

[Ano98e] Anonymous. Index, volume 6, 1998. IEEE Concurrency, 6(4):82–85, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p4082.pdf.

Anonymous:1998:NB

[Ano98f] Anonymous. New books. IEEE Concurrency, 6(1):88, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Anonymous:1998:NNC

[Ano98g] Anonymous. News: NCSA to create support center for distributed applications;and you thought el niño was a problem; Access Para-Scope from Concurrency’s home page. IEEE Concurrency, 6(1):94, January/ March 1998. CODEN IECMFX.

ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1092.pdf.

Anonymous:1999:CCPa

[Ano99a] Anonymous. Calendar/call for papers. IEEE Concurrency, 7 (1):87–88, January/March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1087.pdf.

Anonymous:1999:CCPb

[Ano99b] Anonymous. Calendar/call for papers. IEEE Concurrency, 7(2):88–90, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2088.pdf.

Anonymous:1999:CCPc

[Ano99c] Anonymous. Calendar/call for papers. IEEE Concurrency, 7 (3):95–97, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3095.pdf.

Anonymous:1999:CCPd

[Ano99d] Anonymous. Calendar/call for papers. IEEE Concurrency, 7(4):96–97, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p4096.pdf.

Anonymous:1999:IV

[Ano99e] Anonymous. Index, volume 7. IEEE Concurrency, 7 (4):92–95, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p4092.pdf.

Anonymous:2000:AIV

[Ano00a] Anonymous. Annual index: Volume 8, 2000. IEEE Concurrency, 8(4):94–97, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p4094.pdf.

Anonymous:2000:CCPa

[Ano00b] Anonymous. Calendar/call for papers. IEEE Concurrency, 8 (1):88–89, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p1088.pdf.

Anonymous:2000:CCPb

[Ano00c] Anonymous. Calendar/call for papers. IEEE Concurrency, 8(2):88–90, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2088.pdf.

Anonymous:2000:CCPc

[Ano00d] Anonymous. Calendar/call for papers. IEEE Concurrency, 8(3):85, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3085.pdf.

ATIP:1998:GB

[Asi98] Asian Technology Information Project. Global broadcast: CP-PACS: a look at Japanese supercomputing. IEEE Concurrency, 6(1):89–91, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1089.pdf.

Ben-Ari:1998:EEI

[BAB98] Mordechai Ben-Ari and Alan Burns. Education: Extreme interleavings. IEEE Concurrency, 6(3):90, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3090.pdf.

Bacon:1999:R

[Bac99a] Jean Bacon. 1998 reviewers. IEEE Concurrency, 7 (2):87–88, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2087.pdf.

Bacon:1999:ECC

[Bac99b] Jean Bacon. From the Editorin-Chief: Clusters, open interoperability, and mobility. IEEE Concurrency, 7(2):2–3, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2002.pdf.

Bacon:1999:ECD

[Bac99c] Jean Bacon. From the Editorin-Chief: Defining our scope: The Computer Society’s “systems” magazines. IEEE Concurrency, 7(1):2–3, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p1002.pdf.

Bacon:1999:ECE

[Bac99d] Jean Bacon. From the Editorin-Chief: Evolving concepts for new applications. IEEE Concurrency, 7(3):2–3, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p3002.pdf.

Bacon:1999:EC

[Bac99e] Jean Bacon. From the Editorin-Chief: The new millennium a brave new world? IEEE Concurrency, 7(4):2–3, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1999/pdf/p4002.
pdf.

Bacon:2000:ECD

[Bac00a] Jean Bacon. From the Editorin-Chief: DS Online: Serving the distributed systems community. IEEE Concurrency, 8(2):2–3, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2002.pdf.

Bacon:2000:ECE

[Bac00b] Jean Bacon. From the Editorin-Chief: Expectations and challenges in large-scale distributed systems. IEEE Concurrency, 8(1):2–3, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1002.pdf.

Bacon:2000:ECF

[Bac00c] Jean Bacon. From the Editorin-Chief: The final issue: a time to reflect. IEEE Concurrency, 8 (4):2–3, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4002.pdf.

Bacon:2000:ECG

[Bac00d] Jean Bacon. From the Editorin-Chief: The grand opening of DS Online marks the closing of Concurrency. IEEE Concurrency, 8(3):2–3, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3002.pdf.

Bacon:2000:SRa

[Bac00e] Jean Bacon. Salute to 1999 reviewers. IEEE Concurrency, 8 (1):87–88, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1087.pdf.

Bacon:2000:SRb

[Bac00f] Jean Bacon. Salute to 2000 reviewers. IEEE Concurrency, 8 (4):3, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4003.pdf.

Baentsch:1999:JHR

[BBE^{+}99] Michael Baentsch, Peter Buhler, Thomas Eirich, Frank Höring, and Marcus Oestreicher. JavaCard — from hype to reality. IEEE Concurrency, 7 (4):36–43, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4036.pdf;   http:
//www.computer.org/concurrency/
pd1999/p4036abs.htm.

Bergen:1998:GBT

[Ber98] Shani Bergen. Global broadcast: Thierry van der Pyl and European HPCN trends. IEEE Concurrency, 6(2):11–12, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2011.pdf.

Bestavros:1997:WTR

[Bes97] Azer Bestavros. WWW traffic reduction and load balancing through server-based caching. IEEE Concurrency, 5 (1):56–67, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1056.pdf;   http:
//www.computer.org/concurrency/
pd1997/p1056abs.htm.

Bal:1998:AIT

[BH98] Henri E. Bal and Matthew Haines. Approaches for integrating task and data parallelism. IEEE Concurrency, 6 (3):74–84, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3074.pdf;   http:
//www.computer.org/concurrency/
pd1998/p3074abs.htm.

Bishop:2000:EEI

[Bis00] Matt Bishop. Education: Education in information security. IEEE Concurrency, 8(4):

4–8, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4004.pdf.

Bokhari:1997:BCS

[BN97] Shahid H. Bokhari and David M. Nicol. Balancing contention and synchronization on the Intel Paragon. IEEE Concurrency, 5(2):74–83, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2074.pdf;   http:
//www.computer.org/concurrency/
pd1997/p2074abs.htm.

Beraldi:1999:DST

[BN99] Roberto Beraldi and Libero Nigro. Distributed simulation of timed Petri nets: a modular approach using actors and time warp. IEEE Concurrency, 7(4):52–62, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4052.pdf;   http:
//www.computer.org/concurrency/
pd1999/p4052abs.htm.

Briot:1998:AAA

[Bri98] Jean-Pierre P. Briot. Actors & agents: Agents and concurrent objects: Interview with Les Gasser. IEEE Concurrency, 6 (4):74–77, 81, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4074.pdf.

Burkhart:1997:VRH

[Bur97]  Helmar Burkhart.  Virtual

roundtable: High-performance services à la carte. IEEE Concurrency, 5(3):16–17, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3016b.
pdf.

Bussler:1999:EWW

[Bus99] Christoph Bussler. Enterprisewide workflow management. IEEE Concurrency, 7(3):32–43, July/September 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1999/pdf/p3032.
pdf;   http://www.computer.
org/concurrency/pd1999/p3032abs.
htm.

Callen:1998:ASD

[Cal98] Jeffrey N. Callen. Applications spotlight: Distributed control for unmanned vehicles. IEEE Concurrency, 6(2):16–20, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2016.pdf.

Colajanni:1997:DEP

[CC97] Michele Colajanni and Michele Cermele. DAME: An environment for preserving the efficiency of data-parallel computations on distributed systems. IEEE Concurrency, 5(1):41–55, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1041.pdf;   http:
//www.computer.org/concurrency/
pd1997/p1041abs.htm.

Coutts:1997:MDD

[CE97] Ian A. Coutts and John M. Edwards. Model-driven distributed systems. IEEE Concurrency, 5 (3):55–63, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3055.pdf;   http:
//www.computer.org/concurrency/
pd1997/p3055abs.htm.

Cordsen:1997:VPI

[CGS^{+}97] Jörg Cordsen, Thomas Garnatz, Michael Sander, Anne Gerischer, Marco Dimas Gubitoso, Ute Haack, and Wolfgang Schröder-Preikschat. Vote for Peace: Implementation and performance of a parallel operating system. IEEE Concurrency, 5(2):16–27, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2016.pdf;   http:
//www.computer.org/concurrency/
pd1997/p2016abs.htm.

Charles:1999:FUI

[Cha99] John Charles. Focus: UK Internet spawns anger and innovation. IEEE Concurrency, 7 (3):11–13, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3011.pdf.

Chweh:1997:NPF

[Chw97] Crystal Chweh. New products:

Free C++ template library; templex sees vision of optical RAM; craft-y HPF compiler; getting graphic; real-time vision processing. IEEE Concurrency, 5 (3):81–82, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3081.pdf.

Chweh:1998:NPT

[Chw98] Crystal Chweh. New products:

T3E-1200 outshines its older siblings; data migration facility ports to Irix; Voyager 2.0 blends Corba and Java; virtual flypaper; Labview leads ActiveX lifestyle; small space, big storage; Java performs ScriptEase act. IEEE Concurrency, 6(1):92–93, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1092.pdf.

Chweh:1999:NGD

[Chw99] Crystal Chweh. News: Go distributed. IEEE Concurrency, 7(1):80–82, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1080.pdf.

Clark:1999:NSF

[CKS99] David Clark, Vipin Kumar, and Gil Shif. News: Supercomputing fashion: scalable vectors, the new fall line; mining scientific datasets: Cluster computing: Add processors, save money; new computing complex construction contract [Los Alamos National Laboratory]; IEEE Technical Committee on Parallel Processing; A smarter department of defense; Grid community effort continues. IEEE Concurrency, 7 (4):4–8, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4004.pdf.

Clark:1997:FCB

[Cla97a] David Clark. Focus: Cyberbricks and paradigms. IEEE Concurrency, 5(4):11–12, 17, October/December 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p4011.
pdf.

Clark:1997:NNG

[Cla97b] David Clark. Industry spotlight:

No nukes is good news for supercomputers [nuclear weapons simulation]. IEEE Concurrency, 5(2):11–12, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2011.pdf.

Clark:1997:IBT

[Cla97c] David Clark. Interview: Bluer than blue: how IBM managed it. IEEE Concurrency, 5 (3):8–11, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3008.pdf.

Clark:1998:FAP

[Cla98a] David Clark. Focus: ASCI pathforward: to 30 Tflops and beyond. IEEE Concurrency, 6(2):13–15, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2013.pdf.

Clark:1998:FOP

[Cla98b] David Clark. Focus: OpenMP:

a parallel standard for the masses. IEEE Concurrency, 6 (1):10–12, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1010.pdf.

Clark:1999:SNN

[Cla99] David Clark. Spotlight: Network nirvana and the intelligent device. IEEE Concurrency, 7(2):16–19, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2016.pdf.

Corradi:1999:DLB

[CLZ99] Antonio Corradi, Letizia Leonardi, and Franco Zambonelli. Diffusive load-balancing policies for dynamic applications. IEEE Concurrency, 7(1):22–31, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p1022.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1999/p1022abs.htm.

Chapin:1997:GEI

[CM97] Steve Chapin and Arthur B. (Barney) Maccabe. Guest Editors’ introduction: Multiprocessor operating systems: harnessing the power. IEEE Concurrency, 5(2):14–15, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p2014.pdf.

Cobb:1998:BRT

[Cob98] Maria A. Cobb. Book reviews:

Two books on corba: CORBA Fundamentals and Programming; CORBA Distributed Objects Using Orbix. IEEE Concurrency, 6 (3):86–87, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3085.pdf.

Cano:2000:DBD

[CPSG00] Juan-Carlos Cano, Ana Pont, Julio Sahuquillo, and José-Antonio Gil. The differences between distributed shared memory caching and proxy caching. IEEE Concurrency, 8 (3):45–47, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3045.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3045abs.htm.

Carter:1998:PUA

[CS98] Jeffrey R. Carter and Bo I.

Sanden. Practical use of Ada 95’s concurrency features. IEEE Concurrency, 6(4):47– 56, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4047.pdf;   http:
//www.computer.org/concurrency/
pd1998/p4047abs.htm.

Clark:2000:NBG

[CSFS00] David Clark, Keri Schreiner, Jennifer Ferrero, and Dale Strok. News: Blue Gene and the race toward petaflops capacity; embedded Java development moves ahead; putting teraflops to the test; Corba 3.0 on the way. IEEE Concurrency, 8(1):5–10, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p1005.pdf.

Chweh:1999:NTT

[CST99] Crystal Chweh, John Stankovic, and Dennis Taylor. News: Trafficmaster takes on gridlock; new lab simulates the Internet; Apples are the core of these clusters; growing their own. IEEE Concurrency, 7(2):4–9, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2004.pdf.

dAuriol:1998:BRP

[d’A98] Brian J. d’Auriol. Book review: Parallel Computing Using Optical Interconnections. IEEE Concurrency, 6(3):91, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p3091.pdf.

dAuriol:1999:BRP

[d’A99] Brian J. d’Auriol. Book review: Parallel computing using optical interconnections. IEEE Concurrency, 7(3):91–92, July/ September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1999/pdf/p3091.
pdf.

Dandamudi:1998:SED

[Dan98] Sivarama P. Dandamudi. Sensitivity evaluation of dynamic load sharing in distributed systems. IEEE Concurrency, 6 (3):62–72, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3062.pdf;   http:
//www.computer.org/concurrency/
pd1998/p3062abs.htm.

Dandamudi:1999:RHS

[Dan99] Sivarama P. Dandamudi. Reducing hot-spot contention in shared-memory multiprocessor systems. IEEE Concurrency, 7 (1):48–59, January/March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1048.pdf;   http:
//www.computer.org/concurrency/
pd1999/p1048abs.htm.

Du:1998:ICS

[DCH^{+}98] David H. C. Du, Tai-Sheng Chang, Jenwei Hsieh, Yuewei Wang, and Sangyup Shim. Interface comparisons: SSA versus FC-AL. IEEE Concurrency, 6(2):55–70, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2055.pdf;   http:
//www.computer.org/concurrency/
pd1998/p2055abs.htm.

deDinechin:2000:CEH

[dD00] Christophe de Dinechin. C++ exception handling. IEEE Concurrency, 8(4):72–79, October/December 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p4072.
pdf;   http://www.computer.
org/concurrency/pd2000/p4072abs.
htm.

Decker:1997:VRM

[Dec97] Karsten M. Decker. Virtual roundtable: Matching user requirements. IEEE Concurrency, 5(3):17–19, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3017.pdf.

Devarakonda:1998:GEI

[Dev98] Murthy Devarakonda. Guest Editor’s introduction: The practical aspects of object-oriented programming. IEEE Concurrency, 6 (3):30–33, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3030.pdf.

Devarakonda:2000:GEI

[Dev00] Murthy Devarakonda. Guest Editor’s introduction: Current developments in objectoriented systems. IEEE Concurrency, 8(1):14–15, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1014.pdf.

Decker:1998:ASS

[DJ98] Karsten M. Decker and Mark J.

Johnson. Application specification and software reuse in parallel scientific computing. IEEE Concurrency, 6(2):71–77, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2071.pdf;   http:
//www.computer.org/concurrency/
pd1998/p2071abs.htm.

DiMartino:2000:TPC

[DK00] Beniamino Di Martino and Christoph W. Keßler. Two program comprehension tools for automatic parallelization. IEEE Concurrency, 8(1):37–47, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p1037.
pdf;   http://www.computer.

                                        org/concurrency/pd2000/p1037abs.
                                        htm.

deMeer:1998:TAE

[dRPT98] Hermann de Meer, Jan-Peter Richter, Antonio Puliafito, and Orazio Tomarchio. Tunnel agents for enhanced Internet QoS. IEEE Concurrency, 6(2):30–39, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p2030.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1998/p2030abs.htm.

DiStefano:2000:JKE

[DS00] Antonella Di Stefano and Corrado Santoro. A Java kernel for embedded systems in distributed process control. IEEE Concurrency, 8(4):55–63, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p4055.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p4055abs.htm.

Duncan:1999:BRS

[Dun99a] Michaele E. Duncan. Book review: In Search of Clusters. IEEE Concurrency, 7(2):81, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2081.pdf.

Duncan:1999:BRU

[Dun99b] Ralph Duncan. Book review: Understanding Parallel Supercomputing. IEEE Concurrency, 7 (3):93–94, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3091.pdf.

Eickermann:2000:GBS

[EFH^{+}00] Thomas Eickermann, Wolfgang Frings, Friedel Hossfeld, Stefan Posse, and Gernot Goebbels. Global broadcast: Supercomputerenhanced functional MRI of the human brain. IEEE Concurrency, 8(1):11–13, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1011.pdf.

Elvins:1997:RBV

[Elv97] T. Todd Elvins. Render brokering for volume visualization. IEEE Concurrency, 5 (1):16–25, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1016.pdf;   http:
//www.computer.org/concurrency/
pd1997/p1016abs.htm.

El-Rewini:1997:GEI

[ERH97] Hesham El-Rewini and Wolfang Halang. Guest Editors’ introduction: The engineering of complex distributed computer systems. IEEE Concurrency, 5(4):30–31, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p4030.pdf.

Fairweather:2000:EDL

[FG00a] Peter G. Fairweather and Andrew S. Gibbons. Education: Distributed learning: Two steps forward, one back? one forward, two back? IEEE Concurrency, 8(2):8–9, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2008.pdf.

Felber:2000:POG

[FG00b] Pascal Felber and Rachid Guerraoui. Programming with object groups in CORBA. IEEE Concurrency, 8(1):48–58, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p1048.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p1048abs.htm.

Fleisch:2000:FTC

[FMST00] Brett D. Fleisch, Heiko Michel, Sachin K. Shah, and Oliver E. Theel. Fault tolerance and configurability in DSM coherence protocols. IEEE Concurrency, 8(2):10–21, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2010.pdf;   http:

//www.computer.org/concurrency/
pd2000/p2010abs.htm.

Fekete:1997:VRA

[FP97] Alan Fekete and John M. Potter.

Virtual roundtable: Abstractions in the development of distributed applications. IEEE Concurrency, 5(3):19–20, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3019.pdf.

Garg:1997:MOG

[Gar97]  Vijay K. Garg.  Methods

for observing global properties in distributed systems. IEEE Concurrency, 5(4):69– 77, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4069.pdf;   http:
//www.computer.org/concurrency/
pd1997/p4069abs.htm.

Garraway:1999:PCA

[Gar99]  H. Garraway.  Parallel com-

puter architecture: a hardware/software approach. IEEE Concurrency, 7(2):83–84, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Guessoum:1999:AOA

[GB99] Zahia Guessoum and Jean-Pierre P. Briot. From active objects to autonomous agents. IEEE Concurrency, 7 (3):68–76, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3068.pdf;   http:
//www.computer.org/concurrency/
pd1999/p3068abs.htm.

Green:1997:HTE

[GCSS97] Daron G. Green, Adrian Colbrook, Chris J. Scott, and Mike Surridge. HPCN tools: a European perspective. IEEE Concurrency, 5(3):38–43, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3038.
pdf;   http://www.computer.
org/concurrency/pd1997/p3038abs.
htm.

Guerraoui:1997:GTP

[GGM97] Rachid Guerraoui, Benoît Garbinato, and Karim R. Mazouni. Garf: a tool for programming reliable distributed applications. IEEE Concurrency, 5(4):32– 39, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4032.pdf;   http:
//www.computer.org/concurrency/
pd1997/p4032abs.htm.

Gorton:1997:GEI

[GJ97] Ian Gorton and Innes E. Jelly. Guest Editors introduction: Software engineering for parallel and distributed systems: Challenges and opportunities. IEEE Concurrency, 5(3):12–15, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3012.
pdf.

Gomaa:1997:VRS

[Gom97] Hassan Gomaa. Virtual roundtable:

Software design methods for distributed applications. IEEE Concurrency, 5(3):20, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3020.
pdf.

Giorgi:1997:TFG

[GPPR97] Roberto Giorgi, Cosimo Antonio Prete, Gianpaolo Prina, and Luigi Ricciardi. Trace Factory: Generating workloads for trace-driven simulation of shared-bus multiprocessors. IEEE Concurrency, 5(4):54–68, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4054.pdf;   http:
//www.computer.org/concurrency/
pd1997/p4054abs.htm.

Gilreath:2000:BRJ

[GS00] William F. Gilreath and Benjamin R. Seyfarth. Book reviews: Java Distributed Computing is solid but not flawless; Elements of Network Protocol Design: Network protocol textbook is an asset. IEEE Concurrency, 8(2):86–87, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2086.pdf.

Gill:1998:PCA

[GZ98] Mark A. C. Gill and Albert Y. Zomaya. A parallel collisionavoidance algorithm for robot manipulators. IEEE Concurrency, 6(1):68–78, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1068.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1998/p1068abs.htm.

Hamidzadeh:1998:DSR

[HA98] Babek Hamidzadeh and Yacine Atif. Dynamic scheduling of real-time tasks, by assignment. IEEE Concurrency, 6 (4):14–25, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p4014.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1998/p4014abs.htm.

Hui:1999:ISD

[HC99] Chi-Chung Hui and Samuel T. Chanson. Improved strategies for dynamic load balancing. IEEE Concurrency, 7 (3):58–67, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3058.pdf;   http:
//www.computer.org/concurrency/
pd1999/p3058abs.htm.

Hsieh:1997:HAN

[HDM^{+}97] Jenwei Hsieh, David Hung-Chang C. Du, James A. Mac-Donald, Joseph P. Thomas, and Jack Pugaczewski. HIPPI over ATM networks: Extending connections for distributed computing. IEEE Concurrency, 5(4):40–53, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4040.pdf;   http:
//www.computer.org/concurrency/
pd1997/p4040abs.htm.

Hedberg:1997:DOW

[Hed97a] Sara Reese Hedberg. Democracy online: the White House’s netcasts and town meetings. IEEE Concurrency, 5(1):9–12, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p1009.
pdf.

Hedberg:1997:SRP

[Hed97b] Sara Reese Hedberg. Spotlight: Robots playing soccer? RoboCup poses a new set of challenges in intelligent distributed computing. IEEE Concurrency, 5(4):13–17, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p4013.pdf.

Hedberg:1998:ISC

[Hed98] Sara Reese Hedberg. Industry spotlight: e-Christmas: seeding Europe’s electronic commerce market. IEEE Concurrency, 6(1):13–15, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1013.pdf.

Husemann:1999:OTY

[HH99] Dirk Husemann and Reto Hermann. OpenCard: Talking to your smart card. IEEE Concurrency, 7(3):53–57, July/ September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1999/pdf/p3053.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1999/p3053abs.
                                        htm.

Harizopoulos:2000:HCP

[HHT00] Stavros Harizopoulos, Costas Harizakis, and Peter Triantafillou. Hierarchical caching and prefetching for continuous media servers with smart disks. IEEE Concurrency, 8(3):16–22, July/September 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p3016.
pdf;   http://www.computer.
org/concurrency/pd2000/p3016abs.
htm.

Hwang:1999:DSC

[HJC^{+}99] Kai Hwang, Hai Jin, Edward Chow, Cho-Li Wang, and Zhiwei Xu. Designing SSI clusters with hierarchical checkpointing and single I/O space. IEEE Concurrency, 7(1):60–69, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1060.pdf;   http:
//www.computer.org/concurrency/
pd1999/p1060abs.htm.

Hummel:1997:UPV

[HKS^{+}97] Susan Flynn Hummel, Doug Kimelman, Edith Schonberg, Marsha Tennenhouse, and Dror Zernik. Using program visualization for tuning parallelloop scheduling. IEEE Concurrency, 5(1):26–40, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1026.pdf;   http:
//www.computer.org/concurrency/
pd1997/p1026abs.htm.

Hammer:1997:GB

[HR97] Dieter K. Hammer and Judi Romijn. Global broadcast: Bridging academia and industry — the IPA Workshop on Embedded Systems. IEEE Concurrency, 5(1):89–90, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p1089.pdf.

Heinze:1997:TEP

[HSSO97] Frank Heinze, Lorenz Schäfers, Christian Scheidler, and Wolfgang Obelöer. Trapper: Eliminating performance bottlenecks in a parallel embedded application. IEEE Concurrency, 5 (3):28–37, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p3028.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1997/p3028abs.htm.

Huang:1998:MPC

[HTS98] Yong Gen Huang, Howard Tillotson, and Martin Snaith. Massively parallel computing techniques might improve highway maintenance. IEEE Concurrency, 6(1):58–67, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1058.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1998/p1058abs.htm.

Hundt:2000:HCF

[Hun00] Robert Hundt. HP Caliper: a framework for performance analysis tools. IEEE Concurrency, 8(4):64–71, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4064.pdf;   http:
//www.computer.org/concurrency/
pd2000/p4064abs.htm.

Husemann:1999:SCD

[Hus99] Dirk Husemann. The smart card: Don’t leave home without it. IEEE Concurrency, 7(2):24–27, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2024.pdf;   http:
//www.computer.org/concurrency/
pd1999/p2024abs.htm.

Jajodia:1998:DDI

[Jaj98]  Sushil Jajodia.  Distributed

databases: Interview: Amit Sheth on workflow technology. IEEE Concurrency, 6(2):21–23, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2021.pdf.

Jajodia:1999:DDD

[Jaj99a] Sushil Jajodia. Distributed databases: Data replication gaining popularity. IEEE Concurrency, 7(2):85–86, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2085.pdf.

Jajodia:1999:DDT

[Jaj99b] Sushil Jajodia. Distributed databases: Turing Award winner Jim Gray on the new millenium. IEEE Concurrency, 7(3):9–10, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p3009.pdf.

Jennings:1998:MCS

[JC98] Matthew D. Jennings and Thomas M. Conte. Mobile computing: Subword extensions for video processing on mobile systems. IEEE Concurrency, 6(3):13–16, July/September 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/p3013.
                                        pdf.

Jelly:1997:VRS

[JG97] Innes E. Jelly and Ian Gorton. Virtual roundtable: Software engineering for parallel and distributed systems. IEEE Concurrency, 5(3):16, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p3016a.pdf.

Jovanovic:1999:ORM

[JM99] Milan Jovanovic and Veljko Milutinovic. An overview of reflective memory systems. IEEE Concurrency, 7(2):56–64, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2056.pdf;   http:
//www.computer.org/concurrency/
pd1999/p2056abs.htm.

Jones:1999:GEI

[Jon99]  Alan Jones.  Guest Editor’s

introduction: Mobile computing to go. IEEE Concurrency, 7(2):20–23, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2020.pdf;   http:
//www.computer.org/concurrency/
pd1999/p2020abs.htm.

Joshi:1997:BRT

[Jos97]  Anupam Joshi.  Book re-

views: Two books on distributed computing: Object Oriented Client/Server Internet Environments; Application (Re)Engineering: Building Web-Based Applications and Dealing with Legacies. IEEE Concurrency, 5(4):80–81, October/December 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p4080.
pdf.

Kafil:1998:OTA

[KA98] Muhammad Kafil and Ishfaq Ahmad. Optimal task assignment in heterogeneous distributed computing systems.

IEEE Concurrency, 6(3):42–51, July/September 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/p3042.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1998/p3042abs.
                                        htm.

Kahaner:1997:GB

[Kah97] David Kahaner. Global broadcast: Taiwan’s National Center For High-Performance Computing. IEEE Concurrency, 5(2):91–93, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p2091.pdf.

Kafura:1998:GEI

[KB98] Dennis Kafura and Jean-Pierre Briot. Guest Editors’ introduction: Actors and agents. IEEE Concurrency, 6(2):24–29, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p2024.pdf.

Kon:2000:DMC

[KC00] Fabio Kon and Roy H. Campbell. Dependence management in component-based distributed systems. IEEE Concurrency, 8(1):26–36, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/

pd2000/pdf/p1026.pdf;   http:
//www.computer.org/concurrency/
pd2000/p1026abs.htm.

Kendall:2000:RMA

[Ken00]  Elizabeth A. Kendall.  Role

modeling for agent system analysis, design, and implementation. IEEE Concurrency, 8(2):34–41, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2034.pdf;   http:
//www.computer.org/concurrency/
pd2000/p2034abs.htm.

Kim:1997:PAT

[KK97] Jihong Kim and Yongmin Kim.

Performance analysis and tuning for a single-chip multiprocessor DSP. IEEE Concurrency, 5(1):68–79, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1068.pdf;   http:
//www.computer.org/concurrency/
pd1997/p1068abs.htm.

Kim:1998:FUE

[KP98] Ted H. Kim and Gerald J. Popek.

Frigate: a user-extensible OO file system. IEEE Concurrency, 6(4):26–35, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4026.pdf;   http:

                                        //www.computer.org/concurrency/
                                        pd1998/p4026abs.htm.

Kramer:1997:VRD

[Kra97] Jeffrey Kramer. Virtual roundtable: Distributed software architectures: a structural framework for constructing and reasoning about distributed systems. IEEE Concurrency, 5 (3):21–23, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p3021.pdf.

Krikelis:1997:MAC

[Kri97a] Argy Krikelis. Multimedia — an arena for the culmination of parallel computing. IEEE Concurrency, 5(1):6–8, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p1006.pdf.

Krikelis:1997:MAA

[Kri97b] Argy Krikelis. Multimedia applications: Application-centric parallel multimedia software. IEEE Concurrency, 5(4):78–79, 82, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p4078.pdf.

Krikelis:1997:MAM

[Kri97c] Argy Krikelis. Multimedia applications: Multimedia processing architectures. IEEE Concurrency, 5(3):5–7, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3005.pdf.

Krikelis:1997:MAW

[Kri97d] Argy Krikelis. Multimedia applications: Workshop on Parallel Processing and Multimedia: a new research forum. IEEE Concurrency, 5(2):5–7, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2005.pdf.

Krikelis:1998:MAH

[Kri98a] Argy Krikelis. Multimedia applications: High-performance multimedia applications and the Internet. IEEE Concurrency, 6 (3):17–19, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3017.pdf.

Krikelis:1998:MAM

[Kri98b] Argy Krikelis. Multimedia applications: Managing multimedia information with parallel systems. IEEE Concurrency, 6(2):2–6, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2002.pdf.

Krikelis:1998:MAP

[Kri98c] Argy Krikelis. Multimedia applications: Parallel multimedia servers. IEEE Concurrency, 6(1):16–19, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1016.pdf.

Krikelis:1998:MAS

[Kri98d] Argy Krikelis. Multimedia applications: Scalable multimedia servers. IEEE Concurrency, 6(4):8–10, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p4008.pdf.

Krikelis:1999:MAC

[Kri99a] Argy Krikelis. Multimedia applications: Challenges in delivering multimedia content in mobile environments. IEEE Concurrency, 7(3):77–79, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p3077.pdf.

Krikelis:1999:MAL

[Kri99b] Argy Krikelis. Multimedia applications: Location-dependent multimedia computing. IEEE Concurrency, 7(2):13–15, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2013.pdf.

Krikelis:1999:MAMb

[Kri99c] Argy Krikelis. Multimedia applications: Mobile multimedia considerations. IEEE Concurrency, 7(4):85–87, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4085.pdf.

Krikelis:1999:MAMa

[Kri99d] Argy Krikelis. Multimedia applications: Mobile multimedia: shaping the Infoverse. IEEE Concurrency, 7(1):7–9, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1007.pdf.

Krikelis:2000:MAC

[Kri00a] Argy Krikelis. Multimedia applications: Considerations for a new generation of mobile multimedia communication systems. IEEE Concurrency, 8(2):80–82, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2080.pdf.

Krikelis:2000:MAE

[Kri00b] Argy Krikelis. Multimedia applications: Enhancing visual quality in mobile multimedia. IEEE Concurrency, 8 (3):82–84, July/September 2000.

CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3082.pdf.

Kraemer:1998:CAP

[KS98] Eileen Kraemer and John T. Stasko. Creating an accurate portrayal of concurrent executions. IEEE Concurrency, 6 (1):36–46, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1036.pdf;   http:
//www.computer.org/concurrency/
pd1998/p1036abs.htm.

Karnik:1998:DIM

[KT98] Neeran M. Karnik and Anand R. Tripathi. Design issues in mobile-agent programming systems. IEEE Concurrency, 6 (3):52–61, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3052.pdf;   http:
//www.computer.org/concurrency/
pd1998/p3052abs.htm.

Liszka:1997:PIN

[LAS97] Kathy J. Liszka, John K. Antonio, and Howard Jay Siegel. Problems with interconnection network comparisons: Is an alligator better than an armadillo? IEEE Concurrency, 5(4):18–28, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4018.pdf;   http:
//www.computer.org/concurrency/
pd1997/p4018abs.htm.

Langendoen:1997:MAM

[LBB97] Koen Langendoen, Raoul Bhoedjang, and Henri Bal. Models for asynchronous message handling. IEEE Concurrency, 5(2):28–38, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2028.pdf;   http:
//www.computer.org/concurrency/
pd1997/p2028abs.htm.

Lea:1998:OOA

[Lea98] Doug Lea. Object oriented applications: Patterns and the democratization of concurrent programming. IEEE Concurrency, 6(4):11–13, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4011.pdf.

Levin:1998:BRN

[Lev98] Mikhail P. Levin. Book review:

Numerical Recipes In Fortran 90: The Art Of Parallel Scientific Computing. IEEE Concurrency, 6(4):79, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p4078.pdf.

Lewin:1999:FCC

[Lew99]  David I. Lewin.  Focus: a

changing climate for climate modeling. IEEE Concurrency, 7(2):10–12, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2010.pdf.

Leinberger:1999:GBI

[LK99] William Leinberger and Vipin Kumar. Global broadcast: Information power grid: The new frontier in parallel computing? IEEE Concurrency, 7(4):75–84, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p4075.pdf.

Loques:1998:PRM

[LLV98] Orlando Loques, Julius Leite, and Enrique Vinicio Carrera E. P-RIO: a modular parallelprogramming environment. IEEE Concurrency, 6(1):47–57, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/p1047.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1998/p1047abs.
                                        htm.

Luksch:1997:SSE

[LMR^{+}97] Peter Luksch, Ursula Maier, Sabine Rathmayer, Matthias Weidmann, and Friedemann Unger. Sempa: Software engineering for parallel scientific computing. IEEE Concurrency, 5 (3):64–72, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3064.pdf;   http:
//www.computer.org/concurrency/
pd1997/p3064abs.htm.

Landis:1999:CNS

[LS99] Sean Landis and William Shapiro.

CORBA naming-service evaluation. IEEE Concurrency, 7(4):44–51, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4044.pdf;   http:
//www.computer.org/concurrency/
pd1999/p4044abs.htm.

Lu:2000:ISB

[Lu00] Paul Lu. Implementing scoped behavior for flexible distributed data sharing. IEEE Concurrency, 8(3):63–73, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3063.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3063abs.htm.

Maccabe:1997:BMP

[Mac97] Arthur B. (Barney) Maccabe. Benchmarking multiprocessor OS progress. IEEE Concurrency, 5(2):84–85, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p2084.pdf.

Margenov:1998:BRI

[Mar98] Svetozar Margenov. Book review: An Introduction to Distributed Algorithms. IEEE Concurrency, 6(1):86–87, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Menasce:2000:ERM

[Men00] Daniel A. Menascé. Education: a reference model for designing an e-commerce curriculum. IEEE Concurrency, 8 (1):82–85, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p1082.pdf.

Messerli:1999:PII

[MFGH99] Vincent Messerli, Oscar Figueiredo, Benoit Gennart, and Roger D. Hersch. Parallelizing I/Ointensive image access and processing applications. IEEE Concurrency, 7(2):28–37, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2028.pdf;   http:
//www.computer.org/concurrency/
pd1999/p2028abs.htm.

Minar:2000:HDA

[MGR^{+}00] Nelson Minar, Matthew Gray, Oliver Roup, Raffi Krikorian, and Pattie Maes. Hive: Distributed agents for networking things. IEEE Concurrency, 8(2):24–33, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2024.pdf;   http:
//www.computer.org/concurrency/
pd2000/p2024abs.htm.

Menaud:2000:SEC

[MIB00] Jean-Marc Menaud, Valérie Issarny, and Michel Banâtre. A scalable and efficient cooperative system for Web caches. IEEE Concurrency, 8(3):56–62, July/September 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p3056.
pdf;   http://www.computer.
org/concurrency/pd2000/p3056abs.
htm.

Mikolajczak:1997:DBP

[Mik97] Michael Mikolajczak. Book review: Designing And Building Parallel Programs: Concepts and Tools For Parallel Software Engineering. IEEE Concurrency, 5 (2):88–90, April/June 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Milojicic:1999:TWMa

[Mil99a] Dejan Milojičić. Trend wars: Middleware’s role, today and tomorrow. IEEE Concurrency, 7(2):70–80, April/June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2070.pdf.

Milojicic:1999:TWMb

[Mil99b] Dejan Milojičić. Trend wars: Mobile agent applications. IEEE Concurrency, 7(3):80–90, July/ September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1999/pdf/p3080.
                                        pdf.

Milojicic:1999:TWO

[Mil99c] Dejan Milojičić. Trend wars: Operating systems — now and in the future. IEEE Concurrency, 7(1):12–21, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p1012.pdf.

Milenkovic:2000:AHP

[Mil00a] Aleksandar Milenkovic. Achieving high performance in busbased shared-memory multiprocessors. IEEE Concurrency, 8 (3):36–44, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3036.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3036abs.htm.

Milojicic:2000:GEIa

[Mil00b] Dejan Milojičić. Guest Editor’s introduction: Agent systems and applications. IEEE Concurrency, 8(2):22–23, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2022.pdf.

Milojicic:2000:GEIb

[Mil00c] Dejan Milojičić. Guest Editor’s introduction: Applied operating systems. IEEE Concurrency, 8(4):53–54, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4053.pdf.

Milojicic:2000:TWI

[Mil00d] Dejan Milojičić. Trend wars: Internet technology. IEEE Concurrency, 8(1):70–81, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1070.pdf.

Milojicic:2000:TWS

[Mil00e] Dejan Milojičić. Trend wars:

Security and privacy. IEEE Concurrency, 8(2):70–79, April/ June 2000. CODEN IECMFX.

ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2070.pdf.

Milutinovic:2000:GEI

[Mil00f] Veljko Milutinovic. Guest Editor’s introduction: Caching in distributed systems. IEEE Concurrency, 8(3):14–15, July/ September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p3014.
pdf.

Milojicic:1997:CCS

[MLB^{+}97] Dejan S. Milojičić, Alan Langerman, David L. Black, Michelle Dominijanni, Randall W. Dean, and Steven J. Sears. Concurrency: a case study in remote tasking and distributed IPC in Mach. IEEE Concurrency, 5(2):39–49, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2039.pdf;   http:
//www.computer.org/concurrency/
pd1997/p2039abs.htm.

Milojicic:2000:TWA

[MLS^{+}00] Dejan Milojičić, Brian Lazara, John Stankovic, Wolfgang Schröder Preikschat, Liviu Iftode, Rodger Lea, and Bob Rau. Trend wars: Embedded systems. IEEE Concurrency, 8(4):80–90, October/December 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p4080.
pdf.

Moller:1998:BRD

[Möl98a] Reinhard Möller. Book review:

Distributed Operating Systems: Concepts and Design. IEEE Concurrency, 6(2):93–94, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Moller:1998:BRC

[Möl98b] Reinhard Möller. Book review: On Concurrent Programming. IEEE Concurrency, 6 (1):85–86, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1084.pdf.

Milojicic:2000:TWN

[MPC^{+}00] Dejan Milojičić, Craig Partridge, John Crowcroft, Lance Russell, and Jonathan Smith. Trend wars: Networking. IEEE Concurrency, 8(3):86–96, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3086.pdf.

Mukherjee:2000:WWT

[MRF^{+}00] Shubhendu S. Mukherjee, Steven K.

Reinhardt, Babak Falsafi, Mike Litzkow, Mark D. Hill, David A. Wood, Steven Huss-Lederman, and James R. Larus. Wisconsin Wind Tunnel II: a fast, portable parallel architecture simulator. IEEE Concurrency, 8(4):12–20, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4012.pdf;   http:
//www.computer.org/concurrency/
pd2000/p4012abs.htm.

Mangione-Smith:1998:MCS

[MS98a] William H. Mangione-Smith. Mobile computing and smart spaces. IEEE Concurrency, 6 (4):5–7, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4005.pdf.

Mangione-Smith:1998:MCB

[MS98b] William H. Mangione-Smith. Mobile computing: “Beauty depends on size” — Aristotle. IEEE Concurrency, 6(2):8–10, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2008.pdf.

Mangione-Smith:1999:MCM

[MS99] William H. Mangione-Smith. Mobile computing: MobiCom ’98. IEEE Concurrency, 7 (1):10–11, January/March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1010.pdf.

Pal:1999:BRC

[Pal99] Michelle Pal. Book review: Communication Complexity and Parallel Computing. IEEE Concurrency, 7(2):81–83, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2081.pdf.

Paprzycki:1997:BRI

[Pap97a] Marcin Paprzycki. Book review: An Introduction To Highperformance Scientific Computing. IEEE Concurrency, 5 (3):73–74, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3073.pdf.

Paprzycki:1997:BRP

[Pap97b] Marcin Paprzycki. Book review:

Parallel Computers: Theory and Practice. IEEE Concurrency, 5 (1):81–82, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Paprzycki:2000:BRJ

[Pap00] M. Paprzycki. Book review: Java distributed computing is solid but not flawless. IEEE Concurrency, 8(2):86, April/June 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Passos:1998:BRP

[Pas98] Nelson Luis Passos. Book review:

Pipelined And Parallel Computer Architectures. IEEE Concurrency, 6(2):92, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Passos:1999:BRI

[Pas99]  Nelson Luiz Passos.  Book

review: iWarp — anatomy of a parallel computing system. IEEE Concurrency, 7 (3):92–93, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3091.pdf.

Patnaik:2000:BRH

[Pat00] Lalit M. Patnaik. Book review: High performance cluster computing. IEEE Concurrency, 8(1):86–87, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1086.pdf.

Pernice:1997:BRM

[Per97]  Michael Pernice.  Book re-

view: MPI: The Complete Reference. IEEE Concurrency, 5 (1):80–81, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1080.pdf.

Plale:1998:IAD

[PES^{+}98] Beth Plale, Greg Eisenhauer, Karsten Schwan, Jeremy Heiner, Vernard Martin, and Jeffrey Vetter. From interactive applications to distributed laboratories. IEEE Concurrency, 6(2):78–90, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2078.pdf;   http:
//www.computer.org/concurrency/
pd1998/p2078abs.htm.

Pakin:1997:FME

[PKC97] Scott Pakin, Vijay Karamcheti, and Andrew A. Chien. Fast Messages: Efficient, portable communication for workstation clusters and MPPs. IEEE Concurrency, 5(2):60–73, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p2060.pdf;   http:
//www.computer.org/concurrency/
pd1997/p2060abs.htm.

Pyarali:2000:UPP

[POS^{+}00] Irfan Pyarali, Carlos O’Ryan, Douglas Schmidt, Nanbor Wang, Aniruddha S. Gokhale, and Vishal Kachroo. Using principle patterns to optimize realtime ORBs. IEEE Concurrency, 8(1):16–25, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1016.pdf;   http:
//www.computer.org/concurrency/
pd2000/p1016abs.htm.

Potkonjak:1999:MCL

[Pot99]  Miodrag Potkonjak.  Mobile

computing: Low power and QoS. IEEE Concurrency, 7(4): 9–10, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p4009.pdf.

Price:2000:NCH

[Pri00] Dick Price. News: CAIDA helps train network engineers, 109bit cryptographic key cracked, LosLobos–Linux supercluster. IEEE Concurrency, 8(2):4–7, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd2000/pdf/p2004.
                                        pdf.

Palmer:1998:RCS

[PTT98] Michael E. Palmer, Brian Totty, and Stephen Taylor. Ray casting on shared-memory architectures: Memory-hierarchy considerations in volume rendering. IEEE Concurrency, 6(1):20–35, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1020.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1998/p1020abs.htm.

Parhami:2000:LES

[PYZ00] Behrooz Parhami, Yong Yan, and Xiaodong Zhang. Letter to the Editor: Speedup versus profitup. IEEE Concurrency, 8(1):3–4, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p1003.pdf.  See

[YZ99].

Ranganatban:1997:BRP

[Ran97]  N. Ranganatban.  Book re-

view: Parallel And Distributed Computing Handbook. IEEE Concurrency, 5(3):74–75, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3073.
pdf.

Rehak:2000:NVU

[RBHG00] Daniel R. Rehak, Bill Blackmon, Anne R. Humphreys, and Greg Goth. News: The Virtual University: Developing a dynamic learning management portal; new air traffic control tools enable automated coordination. IEEE Concurrency, 8(3):3–5, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3003.pdf.

Roman:1999:BRL

[RI99] T. Roman and J. S. Institute. Book review: Limits To Parallel Computation: P-Completeness Theory. IEEE Concurrency, 7(1):79, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p1079.pdf.

Russ:1999:HAB

[RRR^{+}99] Samuel H. Russ, Katina Reece, Jonathan Robinson, Brad Meyers, Rajesh Rajan, Laxman Rajagopalan, and Chun-Heong Tan. Hector: An agentbased architecture for dynamic resource management. IEEE Concurrency, 7(2):47–55, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2047.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1999/p2047abs.htm.

Richard:1997:CPR

[RS97] Golden G. Richard III and Mukesh Singhal. Complete Process Recovery: Using vector time to handle multiple failures in distributed systems. IEEE Concurrency, 5(2):50–59, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p2050.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1997/p2050abs.htm.

Rover:1998:STC

[RWMB98] Diane T. Rover, Abdul Waheed, Matt W. Mutka, and Aleksandar M. Bakić. Software tools for complex distributed systems: Toward integrated tool environments. IEEE Concurrency, 6(2):40–54, April/ June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2040.pdf;   http:
//www.computer.org/concurrency/
pd1998/p2040abs.htm.

Sato:1999:GB

[Sat99] Mitsuhisa Sato. Global broadcast: COMPaS: a PC-based SMP cluster. IEEE Concurrency, 7(1):82–86, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1082.pdf.

Schmidt:1997:VRM

[Sch97]  Douglas C. Schmidt.  Vir-

tual roundtable: Meeting the challenges of concurrent software. IEEE Concurrency, 5 (3):23–26, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3023.pdf.

Schmidt:1999:EWS

[Sch99] Marc-Thomas T. Schmidt. The evolution of workflow standards. IEEE Concurrency, 7 (3):44–52, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/

                                        pd1999/pdf/p3044.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1999/p3044abs.htm.

Spalink:2000:MAE

[SHG00] Tammo Spalink, John H. Hartman, and Garth A. Gibson. A mobile agent’s effects on file service. IEEE Concurrency, 8(2):62–69, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2062.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p2062abs.htm.

Skillicorn:1999:SPD

[Ski99] David Skillicorn. Strategies for parallel data mining. IEEE Concurrency, 7(4):26–35, October/December 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1999/pdf/p4026.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1999/p4026abs.
                                        htm.

Sherratt:2000:DDS

[SL00] E. M. Sherratt and C. W. Loftus. Designing distributed services with SDL. IEEE Concurrency, 8(1):59–66, January/ March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p1059.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p1059abs.htm.

Sahuquillo:2000:SDC

[SP00] Julio Sahuquillo and Ana Pont.

Splitting the data cache: a survey. IEEE Concurrency, 8 (3):30–35, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3030.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3030abs.htm.

Silverman:1998:HMA

[SPF98] Harvey F. Silverman, William R.

Patterson III, and James L. Flanagan. The huge microphone array. IEEE Concurrency, 6(4):36–46, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4036.pdf;   http:
//www.computer.org/concurrency/
pd1998/p4036abs.htm.

Silverman:1999:HMA

[SPF99] Harvey F. Silverman, William R.

Patterson III, and James L. Flanagan. The huge microphone array, part 2. IEEE Concurrency, 7(1):32–47, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1032.pdf;   http:
//www.computer.org/concurrency/
pd1999/p1032abs.htm.

Srivastava:1998:MCM

[Sri98] Mani Srivastava. Mobile computing: MobiCom ’97 — beyond the third generation. IEEE Concurrency, 6(1):5–8, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1998/pdf/p1005.pdf.

Small:1998:MCS

[SS98] Christopher Small and Margo I. Seltzer. MiSFIT: Constructing safe extensible systems. IEEE Concurrency, 6(3):34–41, July/September 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/p3034.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd1998/p3034abs.
                                        htm.

Stankovic:1997:VRR

[Sta97]  John A. Stankovic.  Virtual

roundtable: Real-time global virtual machines. IEEE Concurrency, 5(3):26–27, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/p3026.
                                        pdf.

Stachurski:1998:BRV

[Sta98]  Andrzej Stachurski.  Book

review: Verification of Sequential and Concurrent Programs. IEEE Concurrency, 6 (3):88, July/September 1998.

CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3085.pdf.

Stachurski:1999:BRS

[Sta99] Andrzej Stachurski. Book review:

Structured development of parallel programs. IEEE Concurrency, 7(4):88–89, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4088.pdf.

Stpiczynski:1997:BRL

[Stp97]  P. Stpiczynski.  Book re-

view: Local Area Networks: a Client/server Approach. IEEE Concurrency, 5(3):75–76, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p3073.
pdf.

Stpiczynski:1998:BRP

[Stp98] Przemysław Stpiczynski. Book review: Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing. IEEE Concurrency, 6 (2):94, April/June 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Stuben:1997:GBE

[Stü97] Klaus Stüben. Global broadcast: Europort-D: parallel computing for European industry. IEEE Concurrency, 5(4): 7–10, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p4007.pdf.

Sheth:1999:PDN

[SvdAA99] Amit P. Sheth, Wil van der Aalst, and Ismailcem B. Arpinar. Processes driving the networked economy. IEEE Concurrency, 7 (3):18–31, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3018.pdf;   http:
//www.computer.org/concurrency/
pd1999/p3018abs.htm.

Shrivastava:1999:GEI

[SW99] Santosh K. Shrivastava and Stuart M. Wheater. Guest Editors’ introduction: Workflowmanagement systems. IEEE Concurrency, 7(3):16–17, July/ September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1999/pdf/p3016.
pdf.

Szmuc:1997:BRD

[Szm97] Tomasz Szmuc. Book review: A decade of concurrency: reflections and perspectives. IEEE Concurrency, 5(1):82–85, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Szyld:1998:BRP

[Szy98] Daniel B. Szyld. Book review:

Parallel Computation: Models and Methods. IEEE Concurrency, 6(4):79–80, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Taha:1998:BRI

[Tah98]  Thiab R. Taha.  Book re-

view: An Introduction to Parallel Computational Fluid Dynamics. IEEE Concurrency, 6 (4):78, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4078.pdf.

Taylor:1997:WCA

[Tay97] Dennis Taylor. Workstation cluster is all show. IEEE Concurrency, 5(1):13–15, January/ March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p1013.pdf.

Theys:1998:CWA

[TBS98] Mitchell D. Theys, Tracy D.

Braun, and Howard Jay Siegel. Conferences: Widespread acceptance of general-purpose, largescale parallel machines: fact, future, or fantasy? IEEE Concurrency, 6(1):79–83, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1079.pdf.

Thanisch:2000:ACC

[Tha00] Peter Thanisch. Atomic commit in concurrent computing. IEEE Concurrency, 8 (4):34–41, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4034.pdf;   http:
//www.computer.org/concurrency/
pd2000/p4034abs.htm.

Tari:2000:CMC

[THL00] Zahir Tari, Herry Hamidjaja, and Qi Tang Lin. Cache management in CORBA distributed object systems. IEEE Concurrency, 8(3):48–55, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p3048.pdf;   http:
//www.computer.org/concurrency/
pd2000/p3048abs.htm.

Tseng:1998:BMT

[TJD98] Yu-Chee Tseng, Tony Tong-Ying Juang, and Ming-Chih Du. Building a multicasting tree in a high-speed network. IEEE Concurrency, 6 (4):57–67, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4057.pdf;   http:

//www.computer.org/concurrency/
pd1998/p4057abs.htm.

Theilmann:2000:ODM

[TR00] Wolfgang Theilmann and Kurt Rothermel. Optimizing the dissemination of mobile agents for distributed information filtering. IEEE Concurrency, 8(2):53–61, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p2053.pdf;   http:
//www.computer.org/concurrency/
pd2000/p2053abs.htm.

Tripathi:1999:E

[Tri99]  Anand Tripathi.  Education:

Keeping up with the information technology revolution. IEEE Concurrency, 7(3):14–15, July/September 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1999/pdf/p3014.
pdf.

Trobec:1998:BRP

[Tro98] Roman Trobec. Book review:

Parallel Methods for VLSI Layout Design. IEEE Concurrency, 6(1): 87–88, January/March 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Trzaska:1997:BRF

[Trz97] Zdzisław W. Trzaska. Book review: Foundations of Distributed Artificial Intelligence.

IEEE Concurrency, 5(4):82, October/December 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1997/pdf/p4080.
pdf.

Uchihira:1997:HPN

[UHS97] Naoshi Uchihira, Shinichi Honiden, and Toshibumi Seki. Hypersequential programming: a new way to develop concurrent programs. IEEE Concurrency, 5 (3):44–54, July/September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1997/pdf/p3044.pdf;   http:
//www.computer.org/concurrency/
pd1997/p3044abs.htm.

Umar:1998:BRD

[Uma98a] Amjad Umar. Book review: Distributed And Parallel Computing. IEEE Concurrency, 6(4):80–81, October/December 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Umar:1998:BRS

[Uma98b] Amjad Umar. Book review: Second Generation Client/Server Computing. IEEE Concurrency, 6(2):91, April/June 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p2091.pdf.

Vanneschi:1998:GBP

[Van98]  Marco Vanneschi.  Global

broadcast: PQE2000: HPC tools for industrial applications. IEEE Concurrency, 6 (4):68–73, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p4068.pdf.

Vazhenin:1998:BRA

[Vaz98] Alexander Vazhenin. Book review: Advanced Computer Architecture: A Design Space Approach. IEEE Concurrency, 6 (3):87–88, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3085.pdf.

vanderStok:1999:CPM

[vdSAW99] Peter D. V. van der Stok, Erco Argante, and Ian Willers. CoCa: a parallelization model for high-energy physics. IEEE Concurrency, 7(2):38–46, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p2038.pdf;   http:
//www.computer.org/concurrency/
pd1999/p2038abs.htm.

vanderVeen:1998:BRH

[vdV98] Hilda van der Veen. Book review: High Performance Computing: Challenges for Future Systems. IEEE Concurrency, 6 (3):85–86, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3085.pdf.

Vetter:1999:THP

[VS99] Jeffrey Vetter and Karsten Schwan. Techniques for highperformance computational steering. IEEE Concurrency, 7 (4):63–74, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4063.pdf;   http:
//www.computer.org/concurrency/
pd1999/p4063abs.htm.

vanSteen:1999:GWA

[vSHT99] Maarten van Steen, Philip Homburg, and Andrew S. Tanenbaum. Globe: a wide-area distributed system. IEEE Concurrency, 7(1):70–78, January/ March 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p1070.pdf;   http:
//www.computer.org/concurrency/
pd1999/p1070abs.htm.

Watson:2000:RDD

[WA00] Greg Watson and David Abramson. Relative debugging for data-parallel programs: a ZPL case study. IEEE Concurrency, 8(4):42–52, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd2000/pdf/p4042.pdf;   http:
//www.computer.org/concurrency/
pd2000/p4042abs.htm.

Waldo:1998:SRP

[Wal98] Jim Waldo. Spotlight: Remote procedure calls and Java Remote Method Invocation. IEEE Concurrency, 6(3):5–7, July/ September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd1998/pdf/p3005.
pdf.

Wang:1998:OOA

[WC98] Yi-Min Wang and Pi-Yu Emerald Chung. Object-oriented applications: Customization of distributed systems using COM. IEEE Concurrency, 6 (3):8–12, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3008.pdf.

Weissman:1998:BRB

[Wei98] Jon B. Weissman. Book review:

Building Secure and Reliable Network Applications. IEEE Concurrency, 6(1):84–85, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p1084.pdf.

Weissman:2000:EGC

[Wei00]  Jon B. Weissman.  Educa-

tion: Grids in the classroom. IEEE Concurrency, 8 (3):6–9, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p3006.pdf.

Williams:2000:DCP

[WRdS00] Craig Williams, Paul F. Reynolds, Jr., and Bronis R. de Supinski. Delta coherence protocols. IEEE Concurrency, 8 (3):23–29, July/September 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p3023.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p3023abs.htm.

Wojciechowski:2000:NPL

[WS00] Pawel T. Wojciechowski and Peter Sewell. Nomadic Pict: Language and infrastructure design for mobile agents. IEEE Concurrency, 8(2):42–52, April/ June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd2000/pdf/p2042.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd2000/p2042abs.htm.

Weatherford:1999:NGA

[WSFC99] Margaret Weatherford, Keri Schreiner, Jenny Ferrero, and Crystal Chweh. News: Guide for all seasons: Lancaster’s wireless tourist information system; coins (and billions) back Malaysia’s MSC; data encryption: I’d like to solve the puzzle!; click to end world hunger. IEEE Concurrency, 7(3):4–8, July/September 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p3004.pdf.

Wyrzykowski:2000:BRU

[Wyr00] Roman Wyrzykowski. Book review: Using Plapack: Parallel Linear Algebra Package. IEEE Concurrency, 8(3):12–13, July/September 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

URL http://dlib.computer.
org/pd/books/pd2000/pdf/p3012.
pdf.

Yalamov:1998:BRC

[Yal98]  Plamen Y. Yalamov.  Book

review: Concurrent Programming: The Java Programming Language. IEEE Concurrency, 6(3):89, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1998/pdf/p3085.pdf.

Yalamov:1999:BRPa

[Yal99a] Plamen Yalamov. Book review: Parallel Algorithms. IEEE Concurrency, 7(4):89–90, October/December 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1999/pdf/p4088.
                                        pdf.

Yalamov:1999:BRPb

[Yal99b] Plamen Yalamov. Book review: Practical Parallel Programming. IEEE Concurrency, 7 (4):91, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p4088.pdf.

Yalamov:1997:BRT

[YM97] Plamen Y. Yalamov and Svetozar Margenov. Book reviews: Two books on MPI: Parallel Programming with MPI; MPI: The Complete Reference (2nd printing). IEEE Concurrency, 5 (4):81, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p4080.pdf.

Yan:1999:PPE

[YZ99] Yong Yan and Xiaodong Zhang. Performance: Profit-effective parallel computing. IEEE Concurrency, 7(2):65–69, April/ June 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1999/pdf/p2065.pdf. See let-

ter [PYZ00].

Zaki:1999:PDA

[Zak99] Mohammed J. Zaki. Parallel and distributed association mining: a survey. IEEE Concurrency, 7(4):14–25, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

dlib.computer.org/pd/books/
pd1999/pdf/p4014.pdf;   http:
//www.computer.org/concurrency/
pd1999/p4014abs.htm.

Zalewski:1997:BRP

[Zal97a] Janusz Zalewski. Book review:

Pthreads Programming. IEEE Concurrency, 5(1):85–86, January/March 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Zalewski:1997:BRTa

[Zal97b] Janusz Zalewski. Book reviews: Three books on distributed and reactive systems: Time and Probability in Formal Design of Distributed Systems; Interacting Processes: A Multiparty Approach to Coordinated; Temporal Verification of Reactive Systems. IEEE Concurrency, 5(1): 86–87, January/March 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Zalewski:1997:BRTb

[Zal97c] Janusz Zalewski. Book reviews: Two books on distributed real-time systems: Distributed Real-Time Systems: Monitoring, Visualization, Debugging and Analysis, Knowledge-Based Software Development for Real-Time Distributed Systems. IEEE Concurrency, 5(2):86–87, April/ June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1997/pdf/p2086.pdf.

Zomaya:1999:GEI

[ZEGF99] Albert Y. Zomaya, Tarek El-Ghazawi, and Ophir Frieder. Guest Editors’ introduction: Parallel and distributed computing for data mining. IEEE Concurrency, 7(4):11–13, October/December 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1999/pdf/p4011.
                                        pdf.